### v1

# MARS: Reinforcing Multi-Agent Reasoning of LLMs through Self-Play in Strategic Games

Abstract
Developing Large Language Models (LLMs) to cooperate and compete effectively within multi-agent systems is a critical step towards more advanced intelligence. While reinforcement learning (RL) has proven effective for enhancing reasoning in single-agent tasks, its extension to multi-turn, multi-agent scenarios remains underexplored due to the challenges of long-horizon credit assignment and agent-specific advantage estimation. To address these challenges, we introduce MARS, an end-to-end RL framework that incentivizes Multi-Agent Reasoning of LLMs through Self-play in both cooperative and competitive games. MARS features a turn-level advantage estimator that aligns learning signals with each interaction for credit assignment, and an agent-specific advantage normalization to stabilize multi-agent training. By learning with self-play across cooperative and competitive games, the MARS agent trained from Qwen3-4B develops strong strategic abilities that generalize to held-out games with up to 28.728.7% performance improvements. More importantly, the capability acquired through self-play generalizes beyond games, yielding consistent performance gains of multi-agent systems in reasoning benchmarks. When integrated into leading multi-agent systems, our MARS agent achieves significant performance gains of 10.010.0% on AIME and 12.512.5% on GPQA-Diamond. These results establish end-to-end RL training with self-play in strategic games as a powerful approach for developing generalizable multi-agent reasoning capabilities in LLMs. Our code and models are publicly available at https://github.com/thu-nics/MARS.
Figure 1: Evaluation of MARS and two baselines on strategic games and reasoning benchmarks. MARS incentivizes multi-agent reasoning ability through self-play in strategic games and generalizes to improvements of multi-agent systems like MAD and AutoGen on math and QA benchmarks.
1 Introduction
The remarkable capabilities of Large Language Models (LLMs) have revolutionized numerous domains, enabling strong performance on a wide range of tasks from question answering to code generation (Achiam et al., 2023; Team et al., 2023). However, many real-world scenarios, such as negotiation (Bianchi et al., 2024), strategic gameplay (Silver et al., 2016; FAIR et al., 2022), and collaborative software development (Qian et al., 2023; Wu et al., 2024a), inherently involve multiple agents interacting over long horizons. Enabling LLMs to cooperate and compete effectively within multi-agent systems is a critical frontier for advancing artificial intelligence.
While reinforcement learning (RL) has demonstrated remarkable success in enhancing the reasoning capabilities of LLMs (Guo et al., 2025; Team et al., 2025), its extension to multi-turn, multi-agent tasks faces two challenges. First, the problem of long-horizon credit assignment arises in multi-turn interactions. As a sequence of actions collectively leads to the final sparse reward, accurately attributing the contribution of actions in each turn is inherently difficult. Second, multi-agent training couples heterogeneous roles with asymmetric information and different payoff scales, which introduces additional variance in advantage estimation that can destabilize the training process. Recent work (Liu et al., 2025) partially addresses these challenges in adversarial zero-sum settings, but many real-world tasks require both cooperation and competition within multi-agent systems.
In this work, we present MARS, an end-to-end RL framework that incentivizes Multi-Agent Reasoning of LLMs through Self-play in both cooperative and competitive games. We introduce two novel techniques to address the challenges in multi-turn, multi-agent self-play with Group-Relative Policy Optimization (GRPO)  (Shao et al., 2024). First, we propose a simple yet effective turn-level advantage estimator to enable fine-grained credit assignment. This allows the model to accurately attribute long-term outcomes to individual actions and provide learning signals across multiple turns and agents. Second, we propose an agent-specific advantage normalization that stabilizes the training process by calibrating advantage estimates relative to the performance of each agent. This normalization accounts for the heterogeneous roles in multi-agent systems and ensures stable policy updates. By integrating these components, MARS enables robust self-play learning from game outcomes, empowering LLMs to develop strategic abilities and multi-agent reasoning capabilities through cooperation and competition with themselves.
To evaluate the performance and generalization ability of MARS agents, we conduct extensive experiments by training Qwen3-4B in a diverse range of games. Specifically, MARS agents exhibit strong strategic ability in both cooperative and competitive games, with up to 28.728.7% performance improvements in three held-out games. More importantly, the capability acquired through self-play in games further generalizes to improvements of multi-agent systems in reasoning benchmarks. When integrated into both cooperative and competitive multi-agent systems, including AutoGen Wu et al. (2024a) and MAD Liang et al. (2023), our MARS agents achieve performance improvements up to 10.010.0% on AIME and 12.512.5% on GPQA. We further perform ablation studies to validate the effectiveness of our key techniques and analyze the emergent reasoning pattern. The evaluation results establish MARS as a powerful approach for developing generalizable multi-agent reasoning capabilities in LLMs.
In summary, our contributions are as follows:
  • We propose MARS, an end-to-end RL framework that enhances multi-agent reasoning capability through self-play in both cooperative and competitive games.
  • We introduce two novel techniques of turn-level advantage estimation and agent-specific normalization to address the credit assignment and advantage variance in multi-turn, multi-agent RL training for LLMs.
  • We perform extensive experiments and ablation studies to show that MARS incentivize strong strategic ability and multi-agent reasoning capability that are generalizable to held-out games and multi-agent LLM systems.
2 Preliminary
In standard RL fine-tuning for tasks like math solving, the environment is static. The model generates a complete Chain-of-Thought response, and only then is a terminal reward assigned. This process is typically modeled as a token-level Markov Decision Process (MDP). Here, an trajectory consists of a single "turn"—the generation of one complete response with multiple tokens. The state at step $t$, is the sequence of previously generated tokens $(o_{1},...,o_{t-1})$, and the action is the selection of the next token $o_{t}$. Given a task question $q$, the goal is to optimize the token-level policy $\pi(o_{t}|q,o_{<t})$ to produce a sequence that maximizes the final, sparse reward.
In contrast, strategic games introduce a more complex, hierarchical decision-making structure. An entire game, not a single response, constitutes an episode. This is best modeled as a turn-level MDP, where decisions occur at two levels. At the high level, the state $s_{k}$ represents the state of the game at the beginning of turn $k$ (e.g., the board configuration, cards in hand). A high-level action, $a_{k}$, corresponds to the agent’s entire communicative act or move for that turn (e.g., "I place my X in the top-left corner and here is why…"). This action $a_{k}$ is itself a sequence of multiple tokens generated by the LLM’s low-level, autoregressive policy. In this case, the goal is to optimize the turn-level policy $\pi(a_{k}|s_{k})=\prod^{T}_{t=1}\pi(o_{k,t}|s_{k},o_{k,<t})$ to maximize the total return of all turns $R=\sum_{k=1}^{K}r_{k}$.
Figure 2: Overview of MARS. Left column: generating player trajectories through self-play in strategic games. Middle column: naive advantage estimation by GRPO. Right column: advantage estimation by MARS for accurate credit assignment in multi-turn, multi-agent setting.
3 Method
In this section, we introduce MARS, an end-to-end RL framework that enhances multi-agent reasoning ability of LLM through self-play in cooperative and competitive games. We begin by outlining the overall architecture of our training framework, building upon Group-Relative Policy Optimization (GRPO) (Shao et al., 2024). We then detail our primary technical contributions for addressing the credit assignment and advantage estimation challenge in multi-turn, multi-agent training. Finally, we describe the design of our game environments and reward structures. An overview of our proposed method is shown in Fig. 2.
3.1 Self-play with GRPO
To eliminate the extensive computational cost introduce by the critic model in traditional Proximal Policy Optimization (PPO) (Schulman et al., 2017), GRPO rollout each query for multiple times, constructing a group of $G$ responses $\{o^{i}\}_{i=1}^{G}$ and their corresponding rewards $\textbf{r}=\{r^{i}\}_{i=1}^{G}$ for each question, then estimate the advantage of each response by their relative value to their corresponding group, which yields:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o^{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)}\frac{1}{G}\sum^{G}_{i=1}\frac{1}{|o^{i}|}\sum^{|o^{i}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{t}^{i},\varepsilon)$ (1)
where $\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{t}^{i},\varepsilon)=\min\Big[\frac{\pi_{\theta}(o_{t}^{i}|q,o_{<t}^{i})}{\pi_{\theta_{old}}(o_{t}^{i}|q,o_{<t}^{i})}A_{t}^{i},\text{clip}\Big(\frac{\pi_{\theta}(o_{t}^{i}|q,o_{<t}^{i})}{\pi_{\theta_{old}}(o_{t}^{i}|q,o_{<t}^{i})},1-\varepsilon,1+\varepsilon\Big)A_{t}^{i}\Big]$ is the token-level PPO surrogate objective, and the advantages are estimated by $A_{t}^{i}=\frac{r^{i}-\text{mean}(\textbf{r})}{\text{std}(\textbf{r})}$.
In multi-agent self-play, where all players in a strategic game is controlled by the same model, each episode of game results in a multi-turn trajectory for each player, respectively. A naive , we can consider all trajectories from the same game environment as a group, i.e. $\{(s_{k}^{i},a_{k}^{i})_{k=1}^{K^{i}}\}_{i=1}^{G}$ with total return as the terminal reward for each trajectory $\textbf{r}=\{R^{i}\}_{i=1}^{G}$. Then, with an additional summation for multiple turns, GRPO directly generalized to:
$\mathcal{J}_{\text{GRPO}}^{\text{multi}}(\theta)=\mathbb{E}_{s_{k}^{i}\sim P(S),o_{k,t}^{i}\sim\pi_{\theta_{old}}(O|s_{k}^{i},o_{k,<t}^{i})}\frac{1}{G}\sum^{G}_{i=1}\frac{1}{K^{i}}\sum^{K^{i}}_{k=1}\frac{1}{|o^{i}_{k}|}\sum^{|o^{i}_{k}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{k,t}^{i},\varepsilon)$ (2)
where the advantages are estimated by $A_{k,t}^{i}=\frac{R^{i}-\text{mean}(\textbf{r})}{\text{std}(\textbf{r})}$. This assigns equal advantage to all tokens in the multi-turn trajectory, similar to the original GRPO.
3.2 Advantage estimation for multi-turn, multi-agent learning
While Group-Relative Policy Optimization (GRPO) has demonstrated remarkable efficiency and stability in single-turn settings, its direct application to the multi-turn, multi-agent structure of self-play introduces significant challenges of long-horizon credit assignment and agent- specific advantage estimation. To address these issues, we introduce two novel modifications to the GRPO framework, targeting the unique demands of multi-turn and multi-agent learning, respectively.
Turn-level advantage estimator.
To enable fine-grained credit assignment across a long trajectory, we incorporate the sequence of turn-level rewards $\mathbf{r}=\{(r_{k}^{i})_{k=1}^{K}\}_{i=1}^{G}$. This setup is analogous to the "Process Supervision" setting in the original GRPO paper (Shao et al., 2024). In that context, the proposed advantage estimation involves first normalizing all rewards across the entire batch, $\tilde{r}_{k}^{i}=(r_{k}^{i}-\text{mean}(\mathbf{r}))/\text{std}(\mathbf{r})$, and then computing the cumulative sum of these normalized values, $A_{k}^{i}=\sum_{\hat{k}=k}^{K}\tilde{r}_{\hat{k}}^{i}$. While this procedure normalizes the immediate rewards, a critical flaw remains: the resulting advantages $A_{k}^{i}$ are not, in general, zero-mean, which can reintroduce instability into the policy update.
To rectify this, we propose a crucial inversion of these two steps: we first sum, then normalize. Specifically, we begin by calculating the standard Monte Carlo return (or cumulative reward) from each turn $k$ onwards: $R_{k}^{i}=\sum_{\hat{k}=k}^{K}r_{\hat{k}}^{i}$. We then compute the advantage by normalizing these returns to their mean: $A_{k,t}^{i}=R_{k}^{i}-\text{mean}(\textbf{R})$, where $\mathbf{R}$ is the collection of all cumulative rewards in the group. This formulation is equivalent to Generalized Advantage Estimation (GAE) (Schulman et al., 2015) with $\gamma=1$ and $\lambda=1$, where the state-value function $V(s_{k})$ is approximated by a simple, yet effective, baseline—the empirical mean of the batch returns $\mathbb{E}[\mathbf{R}]$. By ensuring the final advantages are properly centered, this method provides a much more stable learning signal for multi-turn decision-making.
Agent-specific advantage normalization.
In many games, the expected return can be highly dependent on a player’s role (e.g., player 0 vs. player 1, or different roles in a cooperative game). Normalizing advantages across a heterogeneous batch of roles would force all players toward a shared baseline, which is statistically inappropriate and can wash out role-specific learning signals.
To address this, we refine our advantage calculation further. We partition the batch of trajectories into sub-groups based on player role and apply the turn-level advantage estimator described above independently within each sub-group $G^{p}$, where $p$ denotes the player index. This ensures that the advantage for a given action is calculated relative to the average outcome for that specific role, providing a more accurate and stable credit assignment in multi-agent settings. This agent-specific, sum-then-normalize advantage estimation constitutes the final objective for MARS:
$\mathcal{J}_{\text{MARS}}$ $(\theta)=\mathbb{E}_{s_{k}^{p,i}\sim P(S),o_{k,t}^{p,i}\sim\pi_{\theta_{old}}(O|s_{k}^{p,i},o_{k,<t}^{p,i})}$ (3)
$\frac{1}{P}\sum^{P}_{p=1}\frac{1}{G_{p}}\sum^{G_{p}}_{i=1}\frac{1}{K^{i}}\sum^{K^{i}}_{k=1}\frac{1}{|o^{p,i}_{k}|}\sum^{|o^{p,i}_{k}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{k,t}^{p,i},\varepsilon)$
where $A_{k,t}^{p,i}=R_{k}^{p,i}-\text{mean}(\textbf{R}^{p})$, $\textbf{R}^{p}$ denotes the total set of cumulative rewards from subgroup $G^{p}$.
3.3 Game environments
For training environments of MARS, we design a portfolio of six strategic, two-player games to cultivate a mixed set of multi-agent capabilities. The portfolio is partitioned into a training set and a more complex, held-out testing set for out-of-distribution (OOD) evaluations. Each pair is designed for a distinct and complementary competency:
  • Perfect-Information Adversarial Games: To build deterministic planning and spatial reasoning, we train on Tic-Tac-Toe. For evaluation, Connect Four serves as a more complex out-of-distribution test due to its vastly larger state space.
  • Imperfect-Information Adversarial Games: To foster reasoning under uncertainty, we train on Kuhn Poker, a simplified variant of poker that teaches probabilistic inference and bluffing. We then test generalization on Leduc Hold’em, which requires a more sophisticated application of these skills.
  • Imperfect-Information Cooperative Games: The inclusion of cooperative games is a key differentiator of our work. Hanabi is essential for developing social intelligence skills like theory of mind and intent recognition. We train on Mini Hanabi and evaluate on the more challenging Simple Hanabi to test the generalization of these collaborative strategies.
Collectively, this diverse portfolio of both adversarial and cooperative challenges ensures the agent develops a robust multi-agent skillset and avoids overfitting to the narrow, adversarial-only mindset of prior work.
3.4 Reward design
While our primary learning signal is derived from unsupervised game outcomes to minimize reward engineering, we incorporate two auxiliary rewards to ensure stable training. The final reward signal is composed of three components:
  • Intrinsic Game Rewards: The primary signal is the intrinsic game outcome. This is a +/-1 reward for a win/loss/draw in Tic-Tac-Toe, chips won or lost in Kuhn Poker (max of 2), and a shared +1 reward per successfully played card in Mini Hanabi (max of 4). To balance these varying scales during multi-game training, we normalize the maximum reward across games to 4 by scaling the Tic-Tac-Toe reward with a factor of 2.
  • Action Format Regularization: To ensure parsable outputs, we provide a small positive reward (+0.05) for validly formatted actions and a large negative penalty (-10.0) that terminates the game for invalid ones, similar to the approach in DeepSeek-R1 (Guo et al., 2025).
  • Response Length Penalty: To encourage conciseness, we apply a turn-level penalty for verbosity, inspired by Kimi k1.5 (Team et al., 2025), which scales linearly for responses longer than a set threshold. The penalty is calculated as:
    $r_{\text{length}}(l)=\alpha\cdot\max\left(0,1-\frac{l-l_{\text{min}}}{l_{\text{max}}-l_{\text{min}}}\right)$ (4)
    where we set $l_{\text{min}}=11$, $l_{\text{max}}=2048$, and the scaling coefficient $\alpha=0.5$.
4 Experiments
In this section, we present a series of extensive experiments designed to validate MARS’s ability to cultivate generalizable multi-agent skills.
Figure 3: Average normalized game returns. Specialist agents not only master their training domains but also generalize effectively to their more complex, held-out counterparts (e.g., from Tic-Tac-Toe to Connect Four). The generalist model achieves consistently high performance across the entire suite of games, establishing it as the most robust and versatile agent.
4.1 Experimental setup
We use Qwen-4B (Yang et al., 2025), a state-of-the-art instruction-tuned model, as our foundation to measure the incremental multi-agent capabilities unlocked by MARS. We train two model types: specialist agents on single games (Tic-Tac-Toe, Kuhn Poker, Mini Hanabi) and a generalist agent on all three simultaneously. Our primary baseline is SPIRAL, a recent work focused on self-play in purely adversarial, zero-sum games (Liu et al., 2025). Our evaluation is structured in four stages:
  1. 1.
    Strategic ability: We first evaluate strategic competence by benchmarking agents against strong Monte Carlo Tree Search (MCTS) or Nash Equilibrium (NE) opponents. Performance is measured over 1000 games on both the training games and a suite of more complex, held-out variants. For the cooperative Hanabi games, we report the standard self-play score.
  2. 2.
    Generalization to multi-agent systems: This is the primary test of this work. We integrate MARS agents into two established systems: the competitive MAD framework (Liang et al., 2023) and the cooperative AutoGen framework (Wu et al., 2024a), measuring performance on downstream math and QA benchmarks.
  3. 3.
    Pattern analysis: We conduct a qualitative analysis of the agent’s chain-of-thought to understand the emergent cognitive skills that MARS develops.
  4. 4.
    Ablation studies: Finally, we perform ablation studies to validate our key algorithmic designs, particularly our novel advantage estimation technique.
Figure 4: Eval curves of Tic-Tac-Toe specialist in Tic-Tac-Toe and Mini Hanabi.
4.2 Strategic ability
Our analysis begins with strategic ability on games, with normalized game return detailed in Fig 3. The MARS framework proves highly effective: specialist agents not only master their training domains and outperform baselines, but also generalize effectively to more complex, held-out counterparts (e.g., from Tic-Tac-Toe to Connect Four). We also observe evidence of cross-category skill transfer. For example, the Tic-Tac-Toe specialist model demonstrates smooth improvement not only on Tic-Tac-Toe, but also the OOD Mini Hanabi (Fig. 4), suggesting MARS cultivates foundational skills like turn-based planning that are broadly beneficial.
Crucially, the generalist model, trained on all environments simultaneously, achieves high performance across the entire suite of games, achieving 28.7% improvement on Leduc Hold’em and 22.9% on Simple Hanabi. Its broad competence across adversarial, cooperative settings establishes it as our most robust and versatile agent overall.
Table 1: Evaluation results on downstream reasoning benchmarks within multi-agent systems. Adversarial-trained agents excel in the competitive MAD framework, while the cooperative-trained agent excels in the collaborative AutoGen framework. The generalist model performs robustly across both. Highest scores are in green; second-highest in light-green, respectively.
Setting Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
single agent Qwen3-4B 60.74 87.60 94.60 39.80 36.70 70.00 57.10 39.39
SPIRAL 63.75 87.50 94.80 51.20 36.70 80.00 58.70 37.37
MARS (Tic.) 63.54 89.10 95.20 46.50 40.00 77.50 57.60 38.89
MARS (Kuhn.) 61.38 87.80 94.50 48.40 33.30 72.50 59.30 33.84
MARS (Hanabi) 62.05 88.10 94.70 48.00 43.30 65.00 58.90 36.36
MARS (Generalist) 62.79 89.90 94.60 52.00 33.30 75.00 59.90 34.85
MAD + Qwen3-4B 72.45 90.20 95.91 80.71 40.00 75.00 87.42 37.88
SPIRAL 73.41 91.60 95.45 81.89 40.00 77.50 87.01 40.40
MARS (Tic.) 75.01 92.20 96.06 83.07 43.33 82.50 86.76 41.12
MARS (Kuhn.) 74.54 91.60 96.21 82.68 40.00 82.50 87.39 41.41
MARS (Hanabi) 73.70 91.40 95.60 82.68 43.33 77.50 87.04 38.38
MARS (Generalist) 75.96 92.80 95.60 83.86 46.67 80.00 87.36 45.45
AutoGen + Qwen3-4B 79.14 93.40 94.69 85.04 56.67 87.50 89.21 47.47
SPIRAL 80.05 94.20 94.47 86.61 60.00 87.50 91.60 45.96
MARS (Tic.) 80.15 94.40 94.69 87.01 60.00 90.00 89.53 45.45
MARS (Kuhn.) 81.54 95.80 94.39 86.61 63.33 92.50 89.65 48.48
MARS (Hanabi) 81.54 94.40 94.54 86.22 66.67 95.00 88.98 44.95
MARS (Generalist) 82.15 95.20 94.54 86.61 66.67 92.50 89.53 50.00
4.3 Generalization to multi-agent systems
The ultimate test of MARS is whether skills honed in games transfer to practical, out-of-domain challenges. We evaluate this on a suite of demanding mathematics and general QA benchmarks, including MATH500 (Cobbe et al., 2021), GSM8K, AQUA-RAT (Ling et al., 2017), AIME24, AMC23, MMLU-STEM (Hendrycks et al., 2020), and GPQA-Diamond (Rein et al., 2024). As a preliminary step, we investigate how our multi-agent curriculum enhances foundational reasoning in a standard single-agent setting. Notably, MARS models achieve significant improvements over both the Qwen-4B baseline on a number of math benchmarks, on par with SPIRAL (Table 1).
We further embed our agents into established Multi-Agent Systems (MAS) to directly measure their collaborative and competitive capabilities. In the competitive MAD debate framework, agents trained on adversarial games show a clear advantage. Conversely, in the collaborative AutoGen framework, the agents trained for cooperation—the Hanabi specialist and the generalist—excel across numerous benchmarks. In particular, the generalist model achieves a striking gain of 10% on AIME, and 12.53% on GPQA-Diamond.
These results provide compelling evidence that MARS forges distinct, transferable skills for both competition and collaboration. These capabilities directly translate to improved performance in downstream multi-agent applications, with the generalist model proving to be the most robust.
4.4 Pattern analysis
To understand the mechanisms behind MARS’s success, we analyze the agent’s chain-of-thought (‘<think>‘) traces. This qualitative analysis reveals the emergence of specific multi-agent skills that are cultivated in games and later transferred to general-purpose systems, as showcased in Table 2.
Table 2: Qualitative analysis of emergent cognitive patterns.
Skill Manifestation in Game-Play Generalize to Multi-Agent Systems
Role
Understanding
The Tic-Tac-Toe specialist recognizes its role as the second player (O) and adopts a defensive strategy.
<think>
Okay, so I am playing the game of Tic-Tac-Toe as the mark O… As the second-move player, I should prioritize blocking the X marks from forming a horizontal, vertical, or diagonal line…
</think>
The same agent, acting as the "negative" debater in MAD, adapts its strategy to refute the opponent.
<think>
Okay, so I need to figure out the minimum number of… The affirmative side just argued for 4. As the negative side assistant, I should try to find an alternative solution and convince the judge…
</think>
Intent
Recognition
The Hanabi specialist infers the intent behind a teammate’s ambiguous hint.
<think>
Okay, so I need to figure out the next move. Player 0 just revealed my only red card. Wait, but the rank remains unknown. Maybe they want me to play this card to the stack?
</think>
The same agent, acting as a user proxy in AutoGen, infers uncertainty from a collaborator’s missing ‘TERMINATE‘ token.
<think>
Okay, let’s see. The assistant gave the answer 17.5 m/s, but did not end the conversation with ’TERMINATE’. Maybe the assistant is not sure with this answer?
</think>
First, MARS develops a Role-Aware Strategy. In Tic-Tac-Toe, the agent explicitly identifies itself as the "second-move player" and adopts a defensive strategy. This skill transfers directly to the MAD debate framework, where the agent recognizes its role as the "negative side assistant" and adopts a negative stance rather than just solving the problem. More profoundly, MARS cultivates Intent Recognition, a key component of Theory of Mind. In the cooperative Hanabi, the agent learns to interpret a teammate’s actions as communications with hidden intent (e.g., "Maybe they want me to play this card?"). This sophisticated skill transfers to the AutoGen framework, where the agent infers a collaborator’s uncertainty from subtle cues, such as a missing ‘TERMINATE‘ token, instead of treating it as a simple error.
These patterns provide qualitative evidence that MARS goes beyond improving benchmark scores; it reshapes the agent’s reasoning process, equipping it with the cognitive toolkit essential for effective multi-agent interaction.
4.5 Ablation studies
To validate our core algorithmic designs, we conduct three targeted ablation studies: (1) comparing self-play to training against a fixed opponent, and (2) sequentially removing the two key components of our advantage estimation: our "sum-then-normalize" turn-level advantage estimator and agent-specific advantage normalization. Full results are in the Appendix.
Self-play vs. fixed-opponent learning.
Training agents against fixed, expert opponents reveals a intuitive and critical flaw: the agents overfit. As shown in Table 3, these models fail to generalize, suffering a severe performance collapse on games outside their direct training domain. This confirms that the adaptive curriculum of self-play is essential for developing robust policies.
Importance of advantage estimation design.
We then ablate our two-part advantage estimation technique. First, reverting from our fine-grained turn-level advantage estimator to a coarse, trajectory-level reward signal causes a significant performance drop, especially for the Tic-Tac-Toe specialist, which is trained on a long-horizon game. Second, removing agent-specific advantage normalization also degrades performance, particularly in asymmetric games-trained model like the Tic-Tac-Toe specialist and the Kuhn Poker specialist where player experiences are distinct. Conversely, the effect is mild in the symmetric game-trained Hanabi speicialist. Together, these results confirm that both components of our advantage estimation are critical for effective learning in complex multi-agent environments.
Table 3: Average normalized game return of the fixed-opponent training ablation. The models demonstrate poor generalization, overfitting to their static opponents and performing significantly worse on out-of-distribution games. Scores below the standard MARS model are highlighted in red.
Training Games Testing Games
Model Tic. Kuhn. Mini. Con.4 Leduc. Simple.
MARS (Tic.) 75.30/32.10 42.90/-77.80 50.48 30.65/14.85 44.70/-17.48 29.75
+ fix opponent 88.00/41.95 18.60/-31.00 34.93 20.35/ 5.65 30.12/-4.65 12.22
MARS (Kuhn.) 69.85/25.50 53.70/-42.70 44.98 27.60/12.70 52.12/38.46 29.35
+ fix opponent 0.00 / 0.00 47.40/-55.30 0.00 0.00 / 0.00 -32.79/-62.38 0.00
Table 4: Average normalized game return of ablation studies on our advantage estimation technique. Both the "sum-then-normalize" approach and per-role normalization are shown to be critical for performance, as their removal leads to a notable drop in game scores. Scores below the standard MARS model are highlighted in red.
Training Games Testing Games
Model Tic. Kuhn. Mini. Con.4 Leduc. Simple.
MARS (Tic.) 75.30/32.10 42.90/-77.80 50.48 30.65/14.85 44.70/-17.48 29.75
+ w/o turn-level. 74.60/24.15 56.40/-31.90 34.80 26.75/12.30 31.40/ 4.75 19.05
+ w/o agent-specific. 82.70/31.20 35.70/-63.40 44.10 25.40/10.50 34.98/18.62 21.72
MARS (Kuhn.) 69.85/25.50 53.70/-42.70 44.98 27.60/12.70 52.12/38.46 29.35
+ w/o turn-level. 63.35/19.65 83.40/-45.40 41.65 29.60/10.85  10.04 /-11.66 22.98
+ w/o agent-specific. 69.55/24.55 45.60/-48.10 40.18 27.00/10.50  14.65 /-27.47 22.42
MARS (Hanabi) 71.90/7.35 39.30/-67.00 55.55 26.75/5.75 16.82/27.13 33.93
+ w/o turn-level. 67.55/10.60 30.30/-25.60 53.20 25.25/3.05 39.97/15.19 30.68
+ w/o agent-specific. 68.15/13.40 42.90/-65.20 52.50 32.10/5.10 26.03/29.21 32.08
5 Related work
LLM-based multi-agent systems.
To further extend the ability of LLMs with collaborative intelligence, LLM-based multi-agent systems have been proposed for various tasks. AutoGen (Wu et al., 2024a) and CAMEL (Li et al., 2023) designs cooperative agents for general reasoning and question answering. MetaGPT (Hong et al., 2024) and ChatDev (Qian et al., 2023) propose static multi-agent workflows with specialized roles for software development tasks. Multi-Agent Debate (Liang et al., 2023) takes a competitive approach by letting LLMs propose and criticize solutions over multiple rounds for the final answer. These works mainly focus on designing workflows for multi-agent interactions with fixed LLMs. Our work takes a complementary approach by training LLMs to enhance their multi-agent capabilities and the trained models can be integrated with these systems.
Reinforcement learning for LLMs.
Reinforcement learning (RL) has emerged as a prominent approach to enhance LLMs’ ability from instruction following to reasoning. To align LLMs with human values, RL from human feedback (RLHF) (Ouyang et al., 2022) and from AI feedback (RLAIF) (Bai et al., 2022) have been widely used in post-training to steer models towards favorable behaviors. The recent success of reasoning models (Jaech et al., 2024; Guo et al., 2025) further popularize RL from verifiable rewards (RLVR) for incentivizing long chain-of-thought reasoning capability of LLMs and have been applied in mathematical reasoning (Shao et al., 2024), coding (Wei et al., 2025), and tool-use (Jin et al., 2025). However, these works primarily consider single-turn or single agent tasks, while our approach extends RL training to multi-turn, multi-agent scenarios.
Self-play training of LLMs.
The great success of achieving superhuman performance in multi-agent games (Silver et al., 2016; Vinyals et al., 2019; Berner et al., 2019) has spurred the use of self-play to train LLMs by playing with themselves. Some work (Chen et al., 2024; Wu et al., 2024b) combines game-theoretic property of self-play to overcome intransitivity of human preference in LLM alignment. Another line of work (FAIR et al., 2022; Xu et al., 2023; 2025) uses self-play to improve the strategic ability of LLM agents in specific games. The most relevant work to ours is SPAG (Cheng et al., 2024) and SPIRAL (Liu et al., 2025), which also train LLMs via self-play and show generalization to reasoning tasks. However, these work focuses on zero-sum game and show generalization in single-agent evaluation. In contrast, our work consider both cooperative and competitive games and demonstrate generalized improvement in multi-agent system.
6 Discussion
In this work, we introduced MARS, a framework that uses self-play in a diverse suite of cooperative and competitive games to cultivate multi-agent capabilities in LLMs. Our findings demonstrate that skills honed in this curriculum directly transfer to and enhance performance in multi-agent systems, establishing self-play as a scalable, autonomous paradigm for training LLM agents.
Despite its promising results, MARS has several limitations that present avenues for future work. First, our current study is confined to two-player games with relatively constrained action spaces. Scaling this approach to more complex, N-player environments will pose significant challenges in credit assignment and strategic modeling. Moreover, while we demonstrate the transferability of learning in strategic games, a particularly compelling direction is the development of more complex "social sandboxes" beyond classic games, designed to train agents in more practical domains like software engineering or scientific discovery.
Ultimately, our work provides strong evidence that the principles of self-play are a powerful engine for progress, paving the way for the next generation of sophisticated, socially intelligent agents.
Reproducibility statement
All code, model checkpoints, and training scripts required to reproduce the findings of this paper are publicly available at https://github.com/thu-nics/MARS. The repository includes all necessary configurations to replicate our key experiments.
Ethics statement
This work adheres to the ICLR Code of Ethics. The primary ethical consideration of our work is the dual-use nature of enhancing LLM agent capabilities. While our goal is to foster beneficial multi-agent skills like collaboration and competition, the resulting models could potentially be applied to malicious ends. We release our code and models to the research community to encourage further study into the safety and alignment of more capable agents.
Appendix A Use of LLMs
In the preparation of this paper, LLMs were utilized as writing assistants. Their use were limited to improving grammar, clarity, and overall readability. The core research, including the methodology, experiments, and analysis, represents the original work of the authors.
Appendix B Implementation details
Framework and software stack.
Our implementation of MARS is built upon ROLL (Wang et al., 2025), a robust open-source codebase for reinforcement learning from LLM feedback. ROLL’s native support for agentic, multi-turn rollouts provided a strong foundation for our system. To achieve high performance, ROLL leverages vLLM for efficient inference during the rollout phase (Kwon et al., 2023) and is built on Megatron-LM for distributed training (Shoeybi et al., 2019). All game environments were implemented using OpenSpiel (Lanctot et al., 2019), ensuring the correctness and standardization of game logic.
Training settings.
We use Qwen-4B as our base model for all experiments. The training is conducted in a fully online manner, where self-play trajectories are immediately used for policy updates. Specialist agents are trained with a batch size of 128 trajectories, while the generalist agent uses a batch size of 384 (128 from each game). All models were trained for a total of 200 optimization steps. For optimization, we employ the Adam optimizer (Kingma, 2014) with a cosine annealing learning rate schedule, which warms up for 10 steps to a peak of $1\times 10^{-6}$ before decaying.
Generation parameters.
During the self-play rollout phase, text is generated via nucleus sampling. We use a temperature of 0.6, a Top-P of 0.99, and a Top-K of 100. These parameters were selected to encourage a diverse yet coherent exploration of strategies.
Hardware configuration.
All training was conducted on a single machine equipped with 8 NVIDIA H100 GPUs.
Appendix C Evaluation details
Strategic ability.
To rigorously assess game-playing proficiency, we evaluate our agents against a suite of strong, fixed opponents. For adversarial games, this includes:
  • Tic-Tac-Toe and Connect Four: MCTS agents with varying simulation counts (100/1000 for Tic-Tac-Toe, 10/100 for Connect Four) to test against different strengths.
  • Kuhn Poker: The exact Nash Equilibrium (NE) policy.
  • Leduc Hold’em: A NE policy approximated by $5\times 10^{8}$ iterations of Counterfactual Regret Minimization (CFR).
In all adversarial games, models are evaluated as both the first-move and second-move player, measured by the average game return. For the cooperative Hanabi variants, performance is measured by the standard self-play game return. In all game settings, agents are evaluated for 1000 games.
For the fixed-opponent ablation studies, the MCTS agent with 100 simulations is used for Tic-Tac-Toe as the opponent, and the NE policy is used for Kuhn Poker.
Generalization to multi-agent systems.
For downstream task evaluation, we test all models on a suite of mathematics and general QA benchmarks. In the standard single-agent setting, we use the evaluation scripts from Qwen2.5-Math (Yang et al., 2024a; b) for MATH500, GSM8K, AQUA-RAT, AIME24, AMC23, and MMLU-STEM, and lm-evaluation-harness (Gao et al., 2024) for GPQA-Diamond. For the multi-agent evaluation, we utilize the MASLab framework (Ye et al., 2025), which integrates multi-agent LLM systems like MAD and AutoGen, for evaluation of all benchmarks.
Across all benchmark evaluations, we maintain consistent generation parameters for a fair comparison: a temperature of 0.6, top-p of 0.95, and top-k of 20. To balance evaluation efficiency with accuracy, we set the maximum output length to 8,192 tokens.
Appendix D Hyperparameters
Table 5: Hyperparameters
Parameter Value
Model Configuration Max Response Length 4096
Max Sequence Length 32768
Dtype bf16
Training Settings Max Steps 200
Train Batch Size 128
Optimizer Adam
Adam parameters ($\beta_{1}$, $\beta_{2}$) (0.9, 0.95)
Learning Scheduler Cosine Annealing
Learning Rate $1\times 10^{-6}$
Weight Decay 0.05
Warm-up Step 10
Gradient Norm Clip 1.0
RL Settings Sampling Temperature 0.5
PPO Epochs 1
(top P, top k) (0.99, 100)
KL Loss True
KL Loss Coefficient 0.2
Entropy Coefficient 0
Dual Clip Loss ture
PPO Policy Clip 0.2
Lambd 0.95
Gamma 1
Appendix E Game observation and prompt
Tic-Tac-Toe
For Tic-Tac-Toe, we provide the agent with a complete observation of the 3x3 game board. The state of each cell—whether it is empty, occupied by ’X’, or occupied by ’O’—is explicitly provided. The prompt clearly indicates which player’s turn it is (’X’ or ’O’) and presents the current board state, asking the agent to select coordinates for its next move from the available empty cells. For example, the game begins with a prompt that provides the empty 3x3 grid and asks the agent to make the first move (Listing 1).
Listing 1: Prompt for Tic-Tac-Toe.
system_prompt:
You are an AI agent that makes optimal decisions to win in the game of Tic-Tac-Toe.
user_prompt:
GAME RULES:
1. Tic-Tac-Toe is a two-player board game played on a three-by-three grid. The grid is 0-indexed, where (0,0) is the top-left corner and (2,2) is the bottom-right corner.
2. Two players take turns placing their marks X and O in empty cells of the grid.
3. The player who first places three of their marks in a horizontal, vertical, or diagonal line wins.
4. If all cells are filled and no player wins, the game ends in a draw.
PLAYER INFORMATION:
1. Your mark is X. You are competing with another player controlling the mark O.
2. In each of your turns:
a. The game state demonstrates the current board with a three-line text grid, where X and O are the marks of the two players, and _ represents empty cells.
b. You need to chose an action to place your mark in an empty cell, based on the given game state and the history of your decisions.
c. All legal actions for the current turn are provided in the format of ‘<X({row},{column})>‘, where X is your mark, and {row} and {column} are integers indicating the row and column of the cell to place your mark.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen action}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><X(0,0)></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
___
___
___
LEGAL ACTIONS:
<X(0,0)>, <X(0,1)>, <X(0,2)>, <X(1,0)>, <X(1,1)>, <X(1,2)>, <X(2,0)>, <X(2,1)>, <X(2,2)>.
Kuhn Poker
For Kuhn Poker, we provide the agent with the observation consists of its single private card (e.g., Jack, Queen, or King) and the complete history actions of two players. The prompt is structured to provide this context clearly, asking the agent to decide on its next action (pass or bet) based on its private information and the betting history. For example, the initial prompt at the beginning of game provides the agent’s private card and the empty action history, asking for the first move (Listing 2).
Listing 2: Prompt for Kuhn Poker.
system_prompt:
You are an AI agent that makes optimal decisions to win in the game of Kuhn Poker.
user_prompt:
GAME RULES:
1. Kuhn Poker is a two-player card game. The deck includes only three cards: King (K) > Queen (Q) > Jack (J).
2. At the start of each game, both player_0 and player_1 place 1 chip into the pot as a blind ante.
3. Each player is dealt a private card, and the third card is set aside unseen.
4. The two players take turns acting, starting with player_0. A player can choose to:
a. <PASS>: place no additional chips into the pot.
b. <BET>: place 1 additional chip into the pot.
5. If a player chooses to <PASS> after the other players <BET>, the betting player wins the pot.
6. If both players choose to <PASS> or both players choose to <BET>, the player with the higher card wins the pot.
PLAYER INFORMATION:
1. You are player_0. You are competing with player_1.
2. In each of your turns:
a. The game state shows your private card and the betting history.
b. You need to choose an action based on your card and the current game state.
c. All legal actions for the current turn are provided in the format of ‘<PASS>‘ or ‘<BET>‘.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen action}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><PASS></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
1. Blind ante: both player_0 and player_1 place 1 chip into the pot.
2. Deal: your card is Jack (J).
LEGAL ACTIONS:
<PASS>, <BET>.
Mini Hanabi
For Mini Hanabi, a cooperative imperfect information game, we provide the agent with a unique observation. An agent observes the cards held by all other players but remains unaware of its own hand. The observation also includes the number of remaining information and life tokens, the cards successfully played on the board (fireworks), and the contents of the discard pile. The prompt is highly structured, detailing the game rules, player-specific information, and strict response formatting. For training, we use MiniHanabi, a simplified 2-player version with 2 colors, 2 ranks, 3-card hands, and 3 information/life tokens. For out-of-distribution evaluation, we use a more complex variant, SimpleHanabi, which features 3 colors, 2 ranks, 5-card hands, 8 information tokens and 3 life tokens. For example, an initial prompt details the starting setup of the chosen variant, including all visible cards and token counts, before asking the first player to make the first move (Listing 3).
Listing 3: Prompt for Mini Hanabi.
system_prompt:
You are an AI agent that makes optimal decisions to achieve the highest score in the game of Hanabi.
user_prompt:
GAME RULES:
1. Hanabi is a cooperative card game for 2 players, player 0 and player 1.
2. The deck consists of 2 colors: Red(denoted by R), Yellow(denoted by Y), with ranks ranging from 1 to 2. Each color contains 4 cards: three of rank 1, and one of rank 2, for a total of 8 cards.
3. Each player holds 3 cards in hand. Players can observe the hand of the other player, but not their own.
4. There are 3 information tokens and 3 life tokens shared by both players.
5. The objective is to play cards in ascending order of rank, from 1 to 2, to their corresponding color stacks, hence achieving the Fireworks’.
6. The players take turns to take one of the following actions:
a. <Play i‘>: play the i-th card from the players own hand (0-indexed). If the card is sequential to the top card of its corresponding color stack, the move is valid and the card is added to the top of the stack, then both players receive 1 point. Otherwise, a life token is lost.
b. <Discard i‘>: discard the i-th card from the players own hand and gain one information token.
c. <Reveal player +1 color c‘>: spend one information token to reveal all cards of color c in the other players hand.
d. <Reveal player +1 rank r‘>: spend one information token to reveal all cards of rank r in the other players hand.
7. After playing or discarding, the player receives a new card from the deck (if remaining).
8. The game ends when:
a. If all color stacks are completed (i.e., all cards of rank 2 are played to their corresponding color stacks), then both players finish the game with the highest possible total score of 4.
b. If deck is depleted, both players finish the game with a total score which equals the sum of the highest ranks of each color stack.
c. If all life tokens are lost before the above two conditions are met, then both players lose all points they have earned so far, and finish the game with a total score of 0.
PLAYER INFORMATION:
1. You will be playing as the player 0.
2. In each of your turns, you will be provided with the current game state information, including the remaining life tokens and information tokens, the current color stacks, the remaining deck size, the discard pile, the hand of the other player, and the revealed information on your own hand.
3. Known cards are denoted by their color and rank. For example, R2 means a red card of rank 2. 4. The current color stacks are represented by the top card of each color stack. In particular, rank 0 denotes an empty stack. For example, Y0 means the yellow stack is still empty.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen action}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><Discard Card 0></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of all life tokens and end of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
1. There are 3 life tokens and 3 information tokens remaining.
2. The top of the color stacks are: R0 Y0.
3. 2 cards remain in the draw pile.
4. The discard pile currently contains: None.
5. The other players hand:
- Card 0 (Y1): the other player believes it is one of the colors [R, Y] and one of the ranks [1, 2].
- Card 1 (R1): the other player believes it is one of the colors [R, Y] and one of the ranks [1, 2].
- Card 2 (R1): the other player believes it is one of the colors [R, Y] and one of the ranks [1, 2].
6. Your own hand, based on the revealed information:
- Card 0: one of the colors [R, Y] and one of the ranks [1, 2].
- Card 1: one of the colors [R, Y] and one of the ranks [1, 2].
- Card 2: one of the colors [R, Y] and one of the ranks [1, 2].
LEGAL ACTIONS:
<Play card 0>, <Play card 1>, <Play card 2>, <Reveal player +1 color R>, <Reveal player +1 color Y>, <Reveal player +1 rank 1>.
Connect Four
For Connect Four, we provide the agent with a complete observation of the 6x7 game board, showing the positions of all ’X’ and ’O’ marks. The prompt is designed to be comprehensive, initially presenting the rules of the game (connecting four pieces to win, draw conditions). It then informs the agent of its assigned mark (’X’ or ’O’) and the opponent’s mark. For each turn, the prompt presents the current board state and asks the agent to choose a column (0-6) to drop its piece. For example, the game begins with a prompt that provides the empty 6x7 grid, explains the rules and player marks, and asks the first player to choose a column (Listing 4).
Listing 4: Prompt for Connect Four.
system_prompt:
You are an AI agent that makes optimal decisions to win in the game of Connect Four.
user_prompt:
GAME RULES:
1. Connect Four is a two-player board game played on a 6x7 grid. Players take turns dropping their pieces into columns.
2. The goal is to connect four of your pieces horizontally, vertically, or diagonally.
3. Pieces fall to the bottom of the column or stack on top of existing pieces.
4. The first player to connect four pieces wins. If the board fills up without a winner, its a draw.
PLAYER INFORMATION:
1. Your mark is X. You are competing with another player controlling the mark O.
2. In each of your turns:
a. The game state shows the current board as a 6x7 grid.
b. You need to choose a column (0-6) to drop your piece, where 0 denotes the leftmost column, 6 denotes the rightmost column.
c. All legal actions are provided as ‘<X({column})>‘, where X is your mark, and {column} is the column number.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen column}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><X(3)></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
_______
_______
_______
_______
_______
_______
LEGAL ACTIONS:
<X(0)>, <X(1)>, <X(2)>, <X(3)>, <X(4)>, <X(5)>, <X(6)>.
Leduc Hold’em
For Leduc Hold’em, we provide the agent with the observation which contains all necessary information for decision-making. The observation is more complex due to two betting rounds and a public card. In the first round, the agent observes its private card and the betting history. In the second round, after a public card is revealed, the observation is updated to include this board card. The prompt provides the agent with its private card, the current betting history, and the public card if revealed, requiring it to make an action in the context of the evolving game state. For example, the initial prompt at the beginning of game provides the agent with its private card and asks for an action in the first betting round before any actions have been taken (Listing 5).
Listing 5: Prompt for Leduc Hold’em.
system_prompt:
You are an AI agent that makes optimal decisions to win in the game of Leduc Poker.
user_prompt:
GAME RULES:
1. Leduc poker is a two-player card game. The deck includes only six cards: two pairs of King (K), Queen (Q), and Jack (J).
2. At the start of each game, both player_0 and player_1 place 1 chip into the pot as a blind ante.
3. Each player is dealt one private card from the deck, and the remaining cards are set aside unseen.
4. The game has two betting rounds. When the first round ends, one public card from the remaining cards of the deck is revealed to both players.
5. The two players take turns acting in the betting rounds, both starting with player_0. A player can choose to:
a. <FOLD>: stop betting and the other player wins the pot.
b. <CALL>: match the current bet. If no bet has been made in the current round, this is equivalent to checking.
c. <RAISE>: first match the current bet and then add n chips to the bet, where n=2‘ in the first round and n=4‘ in the second round. If no bet has been made in the current round, this is equivalent to betting n chips.
6. A maximum of two <RAISE>s are allowed in each round. Each round ends when both players have acted and their bets are equal.
7. If a player chooses to <FOLD>, the other player wins the pot.
8. If neither player chooses to <FOLD>, the second round ends with a showdown:
a. If a player has a pair (private card = public card), the player wins the pot.
b. If neither player has a pair, the player with the higher card (K > Q > J) wins the pot.
c. If two players have the same card, the players split the pot.
PLAYER INFORMATION:
1. You are player_0. You are competing with player_1.
2. In each of your turns:
a. The game state shows your private card, public card (if revealed), and the betting history.
b. You need to choose an action based on your cards and the current game state.
c. All legal actions for the current turn are provided in the format of ‘<FOLD>‘, ‘<CALL>‘, or ‘<RAISE>‘.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen action}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><CALL></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
1. Blind ante: both player_0 and player_1 place 1 chip into the pot.
2. Deal: your card is J.
LEGAL ACTIONS:
<CALL>, <RAISE>.
Appendix F Full results
Table 6: Full results on game-play return in training games
Model TicTacToe KuhnPoker MiniHanabi
MCTS(100 sims) MCTS(1000 sims)
Qwen3-4B 0.403/-0.629 -0.374/-0.687 -0.148/-0.141 1.200
SPIRAL 0.470/-0.624 -0.297/-0.717 -0.301/-0.149 1.494
MARS (Tic.) 0.506/-0.358 -0.278/-0.379 -0.119/-0.142 2.019
MARS (Kuhn.) 0.397/-0.490 -0.388/-0.538 -0.107/-0.103 1.799
MARS (Hanabi) 0.438/-0.644 -0.350/-0.702 -0.123/-0.130 2.222
MARS (Multi.) 0.544/-0.419 -0.302/-0.450 -0.114/-0.141 2.173
+ fix opponent
MARS (Tic.) 0.760/-0.161 -0.378/-0.162 -0.146/-0.090 1.397
MARS (Kuhn.) -1.000/-1.000 -1.000/-1.000 -0.114/-0.117 0.000
MARS (Hanabi) - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) 0.492/-0.517 -0.292/-0.567 -0.104/-0.091 1.392
MARS (Kuhn.) 0.267/-0.607 -0.426/-0.625 -0.074/-0.106 1.666
MARS (Hanabi) 0.351/-0.788 -0.400/-0.811 -0.133/-0.084 2.128
+ w/o agent-specific advantage normalization
MARS (Tic.) 0.654/-0.376 -0.213/-0.366 -0.127/-0.126 1.764
MARS (Kuhn.) 0.391/-0.509 -0.446/-0.567 -0.116/-0.109 1.607
MARS (Hanabi) 0.363/-0.732 -0.364/-0.764 -0.119/-0.128 2.100
Table 7: Full results on game-play return in testing games
Model ConnectFour LeducHoldem SimpleHanabi
MCTS(10 sims) MCTS(100 sims)
Qwen3-4B -0.462/-0.912 -0.902/-0.998 -0.629/-0.691 0.833
SPIRAL -0.383/-0.833 -0.868/-0.995 -0.870/-0.706 0.836
MARS (Tic.) -0.387/-0.703 -0.858/-0.974 -0.518/-0.702 1.785
MARS (Kuhn.) -0.448/-0.746 -0.874/-0.995 -0.460/-0.327 1.761
MARS (Hanabi) -0.465/-0.885 -0.880/-0.998 -0.736/-0.403 2.036
MARS (Multi.) -0.421/-0.780 -0.871/-0.979 -0.483/-0.487 2.205
+ fix opponent
MARS (Tic.) -0.593/-0.887 -0.939/-0.990 -0.632/-0.616 0.733
MARS (Kuhn.) -1.000/-1.000 -1.000/-1.000 -1.124/-1.003 0.000
MARS (Hanabi) - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) -0.465/-0.754 -0.843/-0.989 -0.622/-0.553 1.143
MARS (Kuhn.) -0.408/-0.783 -0.899/-0.998 -0.789/-0.663 1.379
MARS (Hanabi) -0.495/-0.939 -0.886/-0.998 -0.555/-0.483 1.841
+ w/o agent-specific advantage normalization
MARS (Tic.) -0.492/-0.790 -0.882/-0.983 -0.594/-0.460 1.303
MARS (Kuhn.) -0.460/-0.790 -0.908/-0.992 -0.753/-0.769 1.345
MARS (Hanabi) -0.358/-0.898 -0.902/-0.998 -0.664/-0.389 1.925
Table 8: Full results on general benchmarks in standard single agent setting
Model MATH GSM8K AQUA AIME AMC MMLU GPQA
Qwen3-4B 87.60 94.60 39.80 36.70 70.00 57.10 39.39
SPIRAL 87.50 94.80 51.20 36.70 80.00 58.70 37.37
MARS (Tic.) 89.10 95.20 46.50 40.00 77.50 57.60 38.89
MARS (Kuhn.) 87.80 94.50 48.40 33.30 72.50 59.30 33.84
MARS (Hanabi) 88.10 94.70 48.00 43.30 65.00 58.90 36.36
MARS (Multi.) 89.90 94.60 52.00 33.30 75.00 59.90 34.85
+ fix opponent
MARS (Tic.) 89.10 94.80 53.90 50.00 72.50 63.00 37.88
MARS (Kuhn.) 86.20 94.00 45.70 33.30 72.50 53.80 33.84
MARS (Hanabi) - - - - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) 88.80 94.80 48.80 40.00 72.50 57.70 36.36
MARS (Kuhn.) 88.60 94.00 48.40 40.00 72.50 57.90 34.34
MARS (Hanabi) 87.80 94.50 48.40 36.70 70.00 59.30 36.87
+ w/o agent-specific advantage normalization
MARS (Tic.) 88.20 94.50 53.10 36.70 77.50 58.20 39.39
MARS (Kuhn.) 88.10 94.80 53.90 40.00 75.00 58.70 35.86
MARS (Hanabi) 88.60 93.90 46.10 40.00 75.00 57.00 38.89
Table 9: Full results on general benchmarks using the MAD framework
Model MATH GSM8K AQUA AIME AMC MMLU GPQA
Qwen3-4B 90.20 95.91 80.71 40.00 75.00 87.42 37.88
SPIRAL 91.60 95.45 81.89 40.00 77.50 87.01 40.40
MARS (Tic.) 92.20 96.06 83.07 43.33 82.50 86.76 41.12
MARS (Kuhn.) 91.60 96.21 82.68 40.00 82.50 87.39 41.41
MARS (Hanabi) 91.40 95.60 82.68 43.33 77.50 87.04 38.38
MARS (Multi.) 92.80 95.60 83.86 46.67 80.00 87.36 45.45
+ fix opponent
MARS (Tic.) 94.80 94.54 85.04 53.33 90.00 88.57 42.93
MARS (Kuhn.) 95.00 94.24 86.22 53.33 85.00 88.54 45.45
MARS (Hanabi) - - - - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) 93.80 94.77 85.04 60.00 85.00 88.98 47.98
MARS (Kuhn.) 95.00 94.54 86.22 56.67 90.00 89.05 50.00
MARS (Hanabi) 94.20 94.62 85.43 53.33 90.00 88.51 46.97
+ w/o agent-specific advantage normalization
MARS (Tic.) 94.40 94.62 84.65 56.67 87.50 89.33 46.97
MARS (Kuhn.) 95.00 94.69 84.25 56.67 85.00 89.62 47.47
MARS (Hanabi) 94.20 94.47 87.01 63.33 90.00 88.92 43.94
Table 10: Full results on general benchmarks using the AutoGen framework
Model MATH GSM8K AQUA AIME AMC MMLU GPQA
Qwen3-4B 93.40 94.69 85.04 56.67 87.50 89.21 47.47
SPIRAL 94.20 94.47 86.61 60.00 87.50 91.60 45.96
MARS (Tic.) 94.40 94.69 87.01 60.00 90.00 89.53 45.45
MARS (Kuhn.) 95.80 94.39 86.61 63.33 92.50 89.65 48.48
MARS (Hanabi) 94.40 94.54 86.22 66.67 95.00 88.98 44.95
MARS (Multi.) 95.20 94.54 86.61 66.67 92.50 89.53 50.00
+ fix opponent
MARS (Tic.) 92.40 95.38 83.07 43.33 80.00 87.11 43.94
MARS (Kuhn.) 90.20 95.83 83.46 33.33 72.50 86.69 42.42
MARS (Hanabi) - - - - - - -
+ w/o turn-level advantage estimator
MARS (Tic.) 91.00 95.75 82.28 40.00 72.50 86.34 42.93
MARS (Kuhn.) 91.80 95.45 82.68 36.67 82.50 86.79 41.92
MARS (Hanabi) 91.20 95.91 83.86 40.00 75.00 87.33 41.92
+ w/o agent-specific advantage normalization
MARS (Tic.) 91.20 95.83 83.46 40.00 80.00 87.14 38.58
MARS (Kuhn.) 91.20 95.53 84.65 36.67 75.00 86.63 43.94
MARS (Hanabi) 92.60 96.06 82.68 33.33 80.00 87.11 43.43