### v2

Abstract
Developing large language models (LLMs) to cooperate and compete effectively within multi-agent systems (MASs) is a critical step towards more advanced intelligence. While reinforcement learning (RL) has proven effective for enhancing reasoning in single-agent tasks, its extension to multi-turn, multi-agent scenarios remains underexplored due to the challenges of long-horizon credit assignment and agent-specific advantage estimation. To address these challenges, we introduce MARSHAL, an end-to-end RL framework that incentivizes Multi-Agent Reasoning through Self-play witH strAtegic LLMs in both cooperative and competitive games. MARSHAL features a turn-level advantage estimator that aligns learning signals with each interaction for credit assignment, and an agent-specific advantage normalization to stabilize multi-agent training. By learning with self-play across cooperative and competitive games, MARSHAL agent trained from Qwen3-4B develops strong strategic abilities that generalize to held-out games with up to 28.728.7% performance improvements. More importantly, the capability acquired through self-play generalizes beyond games, yielding consistent performance gains of MASs in reasoning benchmarks. When integrated into leading MASs, our MARSHAL agent achieves significant performance gains of up to 10.010.0% on AIME, 6.66.6% on GPQA-Diamond, and 3.53.5% on average across all benchmarks. These results establish end-to-end RL training with self-play in strategic games as a powerful approach for developing generalizable multi-agent reasoning capabilities in LLMs.
Figure 1: Evaluation of MARSHAL and two baselines on strategic games and reasoning benchmarks. MARSHAL incentivizes multi-agent reasoning ability via self-play in strategic games and generalizes to improvements of multi-agent systems like MAD and AutoGen on math and QA benchmarks.
1 Introduction
The remarkable capabilities of large language models (LLMs) have revolutionized numerous domains, enabling strong performance on a wide range of tasks from question answering to code generation (achiam2023gpt; team2023gemini). However, many real-world scenarios, such as negotiation (bianchi2024well), strategic gameplay (silver2016mastering; meta2022human), and collaborative software development (qian2023chatdev; wu2024autogen), inherently involve multiple agents interacting over long horizons. Enabling LLMs to cooperate and compete effectively within multi-agent systems (MASs) is a critical frontier for advancing artificial intelligence.
While reinforcement learning (RL) has demonstrated remarkable success in enhancing the reasoning capabilities of individual LLMs (guo2025deepseek; team2025kimi), its extension to multi-turn, multi-agent tasks faces two critical challenges. First, the problem of long-horizon credit assignment arises in multi-turn interactions. As a sequence of actions can each result in a immediate reward while collectively leading to the final sparse reward, accurately attributing the contribution of the action in each turn is inherently more complex than standard single-turn settings. Second, multi-agent training couples heterogeneous game roles with asymmetric information and different payoff scales, which introduces variance in advantage estimation, destabilizing the training process.
In this work, we present MARSHAL (Multi-Agent Reasoning through Self-play witH strAtegic LLMs), an end-to-end RL framework that incentivizes multi-agent reasoning skills via self-play in strategic games. We introduce two novel techniques to address the challenges in multi-turn, multi-agent self-play with Group-Relative Policy Optimization (GRPO) (shao2024deepseekmath). First, we propose a simple yet effective turn-level advantage estimator to enable fine-grained credit assignment. This allows the model to accurately attribute long-term outcomes to individual actions and provide learning signals across multiple turns and agents. Second, we propose an agent-specific advantage normalization that stabilizes the training process by calibrating advantage estimates relative to the performance of each agent. This normalization accounts for the heterogeneous roles in multi-agent systems and ensures stable policy updates. By integrating these components, MARSHAL enables robust self-play learning from game outcomes, empowering LLMs to develop strategic abilities and multi-agent reasoning skills through cooperation and competition with themselves.
To evaluate the performance and generalization ability of MARSHAL agents, we conduct extensive experiments by training Qwen3-4B in a diverse range of cooperative and competitive games. Specifically, MARSHAL agents exhibit strong strategic ability across all game environments, with up to 28.728.7% performance improvements in three held-out games. More importantly, the capability acquired through self-play in games further generalizes to improvements of multi-agent systems in reasoning benchmarks. When integrated into both cooperative and competitive multi-agent systems, including AutoGen (wu2024autogen) and MAD (liang2023encouraging), our MARSHAL agents achieve performance improvements up to 10.010.0% on AIME, 6.66.6% on GPQA, and 3.53.5% on average across all benchmarks. We further conduct ablation studies to validate the effectiveness of our algorithmic design, complemented by a comprehensive analysis of reasoning patterns and failure modes to understand the successful generalization. The evaluation results establish MARSHAL as a powerful approach for developing generalizable multi-agent reasoning capabilities in LLMs.
In summary, our contributions are as follows:
  • We propose MARSHAL, an end-to-end RL framework that enhances generalizable multi-agent reasoning capability through self-play in a diverse range of strategic games.
  • We introduce two novel techniques of turn-level advantage estimator and agent-specific normalization to address the credit assignment and advantage variance in multi-turn, multi-agent RL training for LLMs.
  • We perform extensive experiments and ablation studies to show that MARSHAL incentivize strong strategic ability and multi-agent reasoning capability that are generalizable to held-out games and multi-agent LLM systems.
2 Preliminary
In standard RL fine-tuning for tasks like math solving, the environment is static. The model generates a complete Chain-of-Thought response, and only then is a terminal reward assigned. This process is typically modeled as a token-level Markov decision process (MDP). Here, an trajectory consists of a single "turn"—the generation of one complete response with multiple tokens. The state at step $t$, is the sequence of previously generated tokens $(o_{1},...,o_{t-1})$, and the action is the selection of the next token $o_{t}$. Given a task question $q$, the goal is to optimize the token-level policy $\pi(o_{t}|q,o_{<t})$ to produce a sequence that maximizes the final, sparse reward.
In contrast, strategic games introduce a more complex, hierarchical decision-making structure. An entire game, not a single response, constitutes an episode. This is best modeled as a turn-level MDP, where decisions occur at two levels. At the high level, the state $s_{k}$ represents the state of the game at the beginning of turn $k$ (e.g., the board configuration, cards in hand). A high-level action, $a_{k}$, corresponds to the agent’s entire output for that turn (e.g., "I place my X in the top-left corner and here is why…"). This action $a_{k}$ is itself a sequence of multiple tokens generated by the LLM’s low-level, autoregressive policy. In this case, the goal is to optimize the turn-level policy $\pi(a_{k}|s_{k})=\prod^{T}_{t=1}\pi(o_{k,t}|s_{k},o_{k,<t})$ to maximize the total return of all turns $R=\sum_{k=1}^{K}r_{k}$.
Figure 2: Overview of MARSHAL. Left column: generating player trajectories through self-play in strategic games. Middle column: naive advantage estimation by GRPO. Right column: advantage estimation by MARSHAL for accurate credit assignment in multi-turn, multi-agent setting.
3 Method
In this section, we introduce MARSHAL, an end-to-end RL framework that enhances multi-agent reasoning ability of LLM through self-play in a diverse range of cooperative and competitive games. We begin by outlining the overall architecture of our training framework, building upon Group-Relative Policy Optimization (GRPO) (shao2024deepseekmath). We then detail our primary technical contributions for addressing the credit assignment and advantage estimation challenge in multi-turn, multi-agent training. Finally, we describe the selection of our game environments and design of reward structures. An overview of our proposed method is shown in Fig. 2.
3.1 Self-play with GRPO
To eliminate the extensive computational cost introduce by the critic model in classic Proximal Policy Optimization (PPO) (schulman2017proximal), GRPO rollout each query for multiple times, constructing a group of $G$ responses $\{o^{i}\}_{i=1}^{G}$ and their corresponding outcome rewards $\textbf{r}=\{r^{i}\}_{i=1}^{G}$, then estimate the advantage of each response by their relative reward to their corresponding group, which assigns equal advantage to the whole sequence (Fig.2 middle), yielding:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim P(Q),\{o^{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}(O|q)}\frac{1}{G}\sum^{G}_{i=1}\frac{1}{|o^{i}|}\sum^{|o^{i}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{t}^{i},\varepsilon)$ (1)
where $\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{t}^{i},\varepsilon)=\min\Big[\frac{\pi_{\theta}(o_{t}^{i}|q,o_{<t}^{i})}{\pi_{\theta_{old}}(o_{t}^{i}|q,o_{<t}^{i})}A_{t}^{i},\text{clip}\Big(\frac{\pi_{\theta}(o_{t}^{i}|q,o_{<t}^{i})}{\pi_{\theta_{old}}(o_{t}^{i}|q,o_{<t}^{i})},1-\varepsilon,1+\varepsilon\Big)A_{t}^{i}\Big]$ is the token-level PPO surrogate objective, and the advantages are estimated by $A_{t}^{i}=\frac{r^{i}-\text{mean}(\textbf{r})}{\text{std}(\textbf{r})}$.
In multi-agent self-play, where all players in a strategic game is controlled by the same model, each episode of game results in a multi-turn trajectory for each player role, respectively. As a naive generalization of the original GRPO to our multi-turn scenario, we start by considering all trajectories from a game environment as a group of response, i.e. $\{(s_{k}^{i},a_{k}^{i})_{k=1}^{K^{i}}\}_{i=1}^{G}$, with total return as the terminal reward for each trajectory $\textbf{r}=\{R^{i}\}_{i=1}^{G}$. Then, with an additional summation for multiple turns, GRPO directly generalized to:
$\mathcal{J}_{\text{GRPO}}^{\text{multi}}(\theta)=\mathbb{E}_{s_{k}^{i}\sim P(S),o_{k,t}^{i}\sim\pi_{\theta_{old}}(O|s_{k}^{i},o_{k,<t}^{i})}\frac{1}{G}\sum^{G}_{i=1}\frac{1}{K^{i}}\sum^{K^{i}}_{k=1}\frac{1}{|o^{i}_{k}|}\sum^{|o^{i}_{k}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{k,t}^{i},\varepsilon)$ (2)
where the advantages are estimated by $A_{k,t}^{i}=\frac{R^{i}-\text{mean}(\textbf{r})}{\text{std}(\textbf{r})}$. This assigns equal advantage to all tokens in the multi-turn trajectory, similar to the original GRPO. For multi-game training, trajectories for each game are naturally considered separate groups and normalized independently.
3.2 Advantage estimation for multi-turn, multi-agent learning
While Group-Relative Policy Optimization (GRPO) has demonstrated remarkable efficiency and stability in single-turn settings, such direct application to the multi-turn, multi-agent structure of self-play introduces significant challenges of long-horizon credit assignment and agent-specific advantage estimation. To address these issues, we introduce two novel modifications to GRPO.
Turn-level advantage estimator.
To enable fine-grained credit assignment across a long trajectory, we incorporate the sequence of turn-level rewards $\mathbf{r}=\{(r_{k}^{i})_{k=1}^{K}\}_{i=1}^{G}$. This setup is analogous to the "Process Supervision" setting in the original GRPO paper (shao2024deepseekmath). In that context, the proposed advantage estimation involves first normalizing all rewards across the entire batch, $\tilde{r}_{k}^{i}=(r_{k}^{i}-\text{mean}(\mathbf{r}))/\text{std}(\mathbf{r})$, and then computing the cumulative sum of these normalized values, $A_{k}^{i}=\sum_{\hat{k}=k}^{K}\tilde{r}_{\hat{k}}^{i}$. While this procedure normalizes the immediate rewards, a critical flaw remains: the advantages $A_{k}^{i}$ are not, in general, zero-mean, which can reintroduce instability into training.
To rectify this, we propose a crucial inversion of these two steps: we first sum, then normalize (Fig.2 right). Specifically, we begin by calculating the standard Monte Carlo return (or cumulative reward) from each turn $k$ onwards: $R_{k}^{i}=\sum_{\hat{k}=k}^{K}r_{\hat{k}}^{i}$. We then compute the advantage by normalizing these returns to their mean: $A_{k,t}^{i}=R_{k}^{i}-\text{mean}(\textbf{R})$, where $\mathbf{R}$ is the collection of all cumulative rewards in the group. This formulation is equivalent to Generalized Advantage Estimation (GAE) (schulman2015high) with $\gamma=1$ and $\lambda=1$, where the state-value function $V(s_{k})$ is approximated by a simple, yet effective, baseline—the empirical mean of the batch returns $\mathbb{E}[\mathbf{R}]$. By ensuring the final advantages are properly centered, this method provides a much more stable learning signal for multi-turn decision-making.
Agent-specific advantage normalization.
In many games, the expected return can be highly dependent on a player’s role (e.g., player 0 vs. player 1, or different roles in a cooperative game). Normalizing advantages across a heterogeneous batch of roles would force all players toward a shared baseline, which is statistically inappropriate and can wash out role-specific learning signals.
To address this, we refine our advantage calculation further. We partition the batch of trajectories into sub-groups based on player role and apply the turn-level advantage estimator described above independently within each sub-group $G^{p}$ (Fig.2 right), where $p$ denotes the player index. This ensures that the advantage for a given action is calculated relative to the average outcome for that specific role, providing a more accurate and stable credit assignment in multi-agent settings. This agent-specific, sum-then-normalize advantage estimation forms the final objective for MARSHAL:
$\mathcal{J}_{\text{MARSHAL}}$ $(\theta)=\mathbb{E}_{s_{k}^{p,i}\sim P(S),o_{k,t}^{p,i}\sim\pi_{\theta_{old}}(O|s_{k}^{p,i},o_{k,<t}^{p,i})}$ (3)
$\frac{1}{P}\sum^{P}_{p=1}\frac{1}{G_{p}}\sum^{G_{p}}_{i=1}\frac{1}{K^{i}}\sum^{K^{i}}_{k=1}\frac{1}{|o^{p,i}_{k}|}\sum^{|o^{p,i}_{k}|}_{t=1}\mathcal{J}_{\text{surr}}(\pi_{\theta};\pi_{\theta_{old}},A_{k,t}^{p,i},\varepsilon)$
where $A_{k,t}^{p,i}=R_{k}^{p,i}-\text{mean}(\textbf{R}^{p})$, $\textbf{R}^{p}$ denotes the total set of cumulative rewards from subgroup $G^{p}$.
3.3 Game environments
To cultivate complementary multi-agent skills for MARSHAL, we design a portfolio of six strategic, two-player games. These games are partitioned into a training set and a more complex, held-out testing set for out-of-distribution (OOD) evaluation:
  • Perfect-information, competitive games: To build deterministic planning and role adaptation, we train on Tic-Tac-Toe, requiring the agent to recognize its strategic position (e.g., first-mover vs. second-mover) and plan accordingly. For evaluation, Connect Four serves as a more complex out-of-distribution test due to its vastly larger state space.
  • Imperfect-information, competitive games: To foster robust reasoning and decision-making under imperfect-information and uncertainty, we train on Kuhn Poker, a simplified poker variant. We evaluate generalization on the more sophisticated Leduc Hold’em.
  • Imperfect-information, cooperative games: The inclusion of cooperative games is a key differentiator of our work. We use Hanabi to develop social intelligence skills like intent recognition and Theory of Mind. Specifically, we train on Mini Hanabi and evaluate on a more challenging variant (Simple Hanabi) to test the generalization of cooperative strategies.
Collectively, this diverse portfolio of both competitive and cooperative challenges ensures the agent develops a robust multi-agent skillset and avoids overfitting to the narrow, competitive-only mindset of prior work (liu2025spiral).
3.4 Reward design
While our primary learning signal is derived from unsupervised game outcomes to minimize reward engineering, we incorporate two auxiliary rewards to ensure stable training. The final reward signal is composed of three components:
  • Intrinsic game rewards: The primary signal is the intrinsic game outcome. This is a +/-1 reward for a win/loss/draw in Tic-Tac-Toe, chips won or lost in Kuhn Poker (max of 2), and a shared +1 reward per successfully played card in Mini Hanabi (max of 4). To balance these varying scales during multi-game training, we normalize the maximum reward across games to 4 by scaling the Tic-Tac-Toe reward with a factor of 2.
  • Action format regularization: To ensure parsable outputs, we provide a small positive reward (+0.05) for validly formatted actions and a large negative penalty (-10.0) that terminates the game for invalid ones, similar to the approach in DeepSeek-R1 (guo2025deepseek).
  • Response length penalty: To encourage conciseness, we apply a turn-level penalty for verbosity, inspired by Kimi k1.5 (team2025kimi), which scales linearly for responses longer than a set threshold. The penalty is calculated as:
    $r_{\text{length}}(l)=\alpha\cdot\max\left(0,1-\frac{l-l_{\text{min}}}{l_{\text{max}}-l_{\text{min}}}\right)$ (4)
    where we set $l_{\text{min}}=11$, $l_{\text{max}}=2048$, and the scaling coefficient $\alpha=0.5$.
4 Experiments
In this section, we present a series of extensive experiments designed to validate MARSHAL’s ability to obtain generalizable multi-agent skills. Additional results can be found in the Appendix.
Figure 3: Average normalized game returns. Specialist agents not only master their training domains but also generalize effectively to their more complex, held-out counterparts (e.g., from Tic-Tac-Toe to Connect Four). The generalist model achieves consistently high performance across the entire suite of games, establishing it as the most robust and versatile agent.
4.1 Experimental setup
We use Qwen-4B (yang2025qwen3), a state-of-the-art instruction-tuned model, as our foundation to measure the incremental multi-agent capabilities unlocked by MARSHAL. We train two model types: specialist agents on single games (Tic-Tac-Toe, Kuhn Poker, Mini Hanabi) and a generalist agent on all three simultaneously. Our primary baseline is SPIRAL, a recent work focused on self-play in purely competitive, zero-sum games (liu2025spiral). Our evaluation is structured in four stages:
  1. 1.
    Strategic ability: We first evaluate strategic competence by benchmarking agents against strong Monte Carlo Tree Search (MCTS) or Nash Equilibrium (NE) opponents. Performance is measured over 1000 games on both the training games and a suite of more complex, held-out variants. For the cooperative Hanabi games, we report the standard self-play score.
  2. 2.
    Generalization to multi-agent systems: This is the primary test of this work. We integrate MARSHAL agents into two established systems: the competitive MAD framework (liang2023encouraging) and the cooperative AutoGen framework (wu2024autogen), measuring performance on downstream math and QA benchmarks.
  3. 3.
    Pattern analysis: We conduct a qualitative analysis of the agent’s reasoning process and a quantitative analysis of failure modes to understand the successful generalization to MASs.
  4. 4.
    Ablation studies: Finally, we perform ablation studies to validate our key algorithmic designs, particularly our novel advantage estimation technique.
Figure 4: Eval curves of the Tic-Tac-Toe specialist in Tic-Tac-Toe and Mini Hanabi.
4.2 Strategic ability
Our analysis begins with strategic ability on games, with normalized game return detailed in Fig 3. The MARSHAL framework proves highly effective: specialist agents not only master their training domains and outperform baselines, but also generalize effectively to more complex, held-out counterparts (e.g., from Tic-Tac-Toe to Connect Four). We also observe evidence of cross-category skill transfer. For example, the Tic-Tac-Toe specialist model demonstrates smooth improvement not only on Tic-Tac-Toe, but also the OOD Mini Hanabi (Fig. 4), suggesting MARSHAL cultivates foundational skills like turn-based planning that are broadly beneficial.
Crucially, the generalist model, trained on all environments simultaneously, achieves high performance across the entire suite of games, achieving 28.7% improvement on Leduc Hold’em and 22.9% on Simple Hanabi. Its broad competence across competitive and cooperative settings establishes it as our most robust agent overall.
Table 1: Evaluation results on downstream reasoning benchmarks within multi-agent systems. Competitive game-trained agents excel in the competitive MAD framework, while the cooperative-trained agent excels in the cooperative AutoGen framework. The generalist model performs robustly across both. Bold and underlined indicate the best and second-best scores, respectively.
Setting Model Math QA
Average MATH GSM8K AQUA AIME AMC MMLU GPQA
Qwen3-4B 60.74 87.60 94.60 39.80 36.70 70.00 57.10 39.39
SPIRAL 63.75 87.50 94.80 51.20 36.70 80.00 58.70 37.37
MARSHAL
  Tic-Tac-Toe 63.54 89.10 95.20 46.50 40.00 77.50 57.60 38.89
  Kuhn Poker 61.38 87.80 94.50 48.40 33.30 72.50 59.30 33.84
  Mini Hanabi 62.05 88.10 94.70 48.00 43.30 65.00 58.90 36.36
Single Agent Generalist 62.79 89.90 94.60 52.00 33.30 75.00 59.90 34.85
Qwen3-4B 72.45 90.20 95.91 80.71 40.00 75.00 87.42 37.88
SPIRAL 73.41 91.60 95.45 81.89 40.00 77.50 87.01 40.40
MARSHAL
  Tic-Tac-Toe 75.01 92.20 96.06 83.07 43.33 82.50 86.76 41.12
  Kuhn Poker 74.54 91.60 96.21 82.68 40.00 82.50 87.39 41.41
  Mini Hanabi 73.70 91.40 95.60 82.68 43.33 77.50 87.04 38.38
MAD (Compe- titive) Generalist 75.96 92.80 95.60 83.86 46.67 80.00 87.36 45.45
Qwen3-4B 79.14 93.40 94.69 85.04 56.67 87.50 89.21 47.47
SPIRAL 80.05 94.20 94.47 86.61 60.00 87.50 91.60 45.96
MARSHAL
  Tic-Tac-Toe 80.15 94.40 94.69 87.01 60.00 90.00 89.53 45.45
  Kuhn Poker 81.54 95.80 94.39 86.61 63.33 92.50 89.65 48.48
  Mini Hanabi 81.54 94.40 94.54 86.22 66.67 95.00 88.98 44.95
AutoGen (Coope- rative) Generalist 82.15 95.20 94.54 86.61 66.67 92.50 89.53 50.00
4.3 Generalization to multi-agent systems
The ultimate test of MARSHAL is whether skills honed in games transfer to practical, out-of-domain challenges. We evaluate this on a suite of demanding mathematics and general QA benchmarks, including MATH500 (cobbe2021training), GSM8K, AQUA-RAT (ling2017program), AIME24, AMC23, MMLU-STEM (hendrycks2020measuring), and GPQA-Diamond (rein2024gpqa). As a preliminary step, we investigate how our game tasks enhances foundational reasoning in a standard single-agent setting. Notably, MARSHAL models achieve notable improvements over both the Qwen-4B baseline on a number of math benchmarks, on par with SPIRAL (Table 1).
We further embed our agents into established multi-agent systems to directly measure their cooperative and competitive capabilities. In the competitive MAD debate framework, agents trained on competitive games show a clear advantage. Notably, the generalist agent is able to achieve an average of 3.51% over the original Qwen3-4B across all benchmarks. Conversely, in the cooperative AutoGen framework, the agents trained for cooperation—the Hanabi specialist and the generalist—excel across numerous benchmarks. In particular, the generalist model achieves a striking gain of 6.57% on GPQA-Diamond in the MAD framework, and 10.00% on AIME in the AutoGen framework.
These results provide compelling evidence that MARSHAL forges distinct, transferable skills for both competition and collaboration. These capabilities directly translate to improved performance in downstream multi-agent applications, with the generalist model proving to be the most robust.
4.4 Reasoning pattern and failure mode analysis
To understand the mechanisms behind MARSHAL’s successful generalization, we analyze the agent’s reasoning (<think>) traces. This qualitative analysis reveals the emergence of key multi-agent skills cultivated in our mix of strategic games. We highlight two representative patterns corresponding to the competitive and cooperative training environments, as showcased in Table 2.
First, reflecting the skills honed in competitive games, MARSHAL develops a role-aware strategy. In Tic-Tac-Toe, the agent explicitly identifies itself as the "second-move player" and adopts a defensive strategy. This skill transfers directly to the MAD debate framework, where the agent recognizes its role as the "negative side assistant" and adopts a negative stance rather than just solving the problem.
More profoundly, MARSHAL cultivates intent recognition, a key component of Theory of Mind. In the cooperative Hanabi, the agent learns to interpret a teammate’s actions as communications with hidden intent (e.g., "Maybe they want me to play this card?"). This sophisticated skill transfers to the AutoGen framework, where the agent infers a collaborator’s uncertainty from subtle cues, such as a missing TERMINATE token, instead of treating it as a simple error. Together, these patterns provide qualitative evidence that MARSHAL goes beyond improving benchmark scores; it equips the agent with the cognitive toolkit essential for effective multi-agent interaction.
To further substantiate these qualitative observations with quantitative evidence, we performed a failure mode analysis on the GPQA-Diamond benchmark within the MAD framework. Adopting the taxonomy of cemri2025multi, we categorized failures into System Design Issues (e.g., formatting errors, loop repetition), Inter-Agent Misalignment (e.g., multi-agent reasoning failures like ignoring peers or task derailment), and Task Verification issues.
As shown in Fig 5 (left), while MARSHAL improves basic instruction following (reducing System Design Issues by $\sim$7%), the reduction in Inter-Agent Misalignment is significantly larger (11.5%). To understand the drivers of this strategic improvement, we further decompose the Inter-Agent Misalignment category into its sub-categories. The breakdown reveals that the performance gains are primarily driven by reductions in Task Derailment and Ignored Other Agent’s Input. This indicates that MARSHAL agent are actively listening to peers and maintaining focus on its objective, validating our hypothesis that game-theoretic self-play cultivates transferable multi-agent reasoning skills.
Table 2: Qualitative analysis of emergent reasoning patterns.
Skill Manifestation in Game-Play Generalize to Multi-Agent Systems
Role
Understanding
The Tic-Tac-Toe specialist recognizes its role as the second player (O) and adopts a defensive strategy.
<think>
Okay, so I am playing the game of Tic-Tac-Toe as the mark O… As the second-move player, I should prioritize blocking the X marks from forming a horizontal, vertical, or diagonal line…
</think>
The same agent, acting as the "negative" debater in MAD, adapts its strategy to refute the opponent.
<think>
Okay, so I need to figure out the minimum number of… The affirmative side just argued for 4. As the negative side assistant, I should try to find an alternative solution and convince the judge…
</think>
Intent
Recognition
The Hanabi specialist infers the intent behind a teammate’s ambiguous hint.
<think>
Okay, so I need to figure out the next move. Player 0 just revealed my only red card. Wait, but the rank remains unknown. Maybe they want me to play this card to the stack?
</think>
The same agent, acting as a user proxy in AutoGen, infers uncertainty from a collaborator’s missing ‘TERMINATE‘ token.
<think>
Okay, let’s see. The assistant gave the answer 17.5 m/s, but did not end the conversation with ’TERMINATE’. Maybe the assistant is not sure with this answer?
</think>
Figure 5: Percentage of different failure modes in GPQA-Dimond. Through MARSHAL training, the generalist agent significantly reduces the occurrence of Inter-Agent Misalignment.
4.5 Ablation studies
To validate our core algorithmic designs, we conduct two targeted ablation studies: (1) comparing self-play to training against a fixed opponent, and (2) sequentially removing the two key algorithmic components: our turn-level advantage estimator and agent-specific advantage normalization.
Self-play vs. fixed-opponent learning.
Training agents against fixed, expert opponents reveals a intuitive and critical flaw: the agents overfit. As shown in Table 3, these models fail to generalize, exhibiting significant performance drop on games outside their direct training domain. This is particularly evident with the Kuhn Poker specialist, which suffers a severe degradation on non-poker games—a clear case of strategic mode collapse. These results confirm that the adaptive curriculum provided by self-play is essential for the development of robust, generalizable policies.
Figure 6: Return distribution of the generalist MARSHAL agent as different player roles on Tic-Tac-Toe and Mini Hanabi.
Analysis of advantage estimation design.
We then ablate our two-part advantage estimation technique. First, reverting from our fine-grained turn-level advantage estimator to a coarse, trajectory-level reward signal causes a significant performance drop, especially for the Tic-Tac-Toe specialist, which is trained on a long-horizon game. Second, removing agent-specific advantage normalization also degrades performance, particularly in competitive games-trained model like the Tic-Tac-Toe specialist and the Kuhn Poker specialist where player experiences and agent return are distinct (Fig. 6). Conversely, the effect is mild in the Hanabi specialist, which is trained on a cooperative game with similar return distribution between players (Fig. 6). Together, these results confirm that both components of our advantage estimation are critical for effective learning in complex multi-turn, multi-agent environments.
Table 3: Generalization comparison between MARSHAL (self-play) and its fixed-opponent variant. The latter exhibits significant overfitting to static environments and opponents. Values denote average normalized game returns; for competitive games, entries indicate first-move / second-move returns. Underlined scores indicate performance degradation compared to the standard MARSHAL model.
Model Training Games Testing Games
Tic-Tac-Toe Kuhn Poker
Mini
Hanabi
Connect Four
Leduc
Hold’em
Simple
Hanabi
MARSHAL (Tic-Tac-Toe) 75.30 / 32.10 74.15 / 3.42 50.48 30.65 / 14.85 58.36 / 27.65 29.75
w/ fixed opponent 88.00 / 41.95 63.15 / 28.84 34.93 20.35 / 5.65 47.38 / 35.55 12.22
MARSHAL (Kuhn Poker) 69.85 / 25.50 79.04 / 22.49 44.98 27.60 / 12.70 63.94 / 62.10 29.35
w/ fixed opponent 0.00 / 0.00 76.19 / 15.64 0.00 0.00 / 0.00 0.00 / 0.00 0.00
Table 4: Ablation results for algorithmic design. Both our turn-level advantage estimator and agent-specific advantage normalization prove essential for performance. Notation follows Table 3.
Model Training Games Testing Games
Tic-Tac-Toe Kuhn Poker
Mini
Hanabi
Connect Four
Leduc
Hold’em
Simple
Hanabi
MARSHAL (Tic-Tac-Toe) 75.30 / 32.10 74.15 / 3.42 50.48 30.65 / 14.85 58.36 / 27.65 29.75
w/o Turn-Level. 74.60 / 24.15 80.26 / 28.35 34.80 26.75 / 12.30 48.34 / 41.34 19.05
w/o Agent-Specific. 82.70 / 31.20 70.89 / 11.24 44.10 25.40 / 10.50 51.04 / 49.88 21.72
MARSHAL (Kuhn Poker) 69.85 / 25.50 79.04 / 22.49 44.98 27.60 / 12.70 63.94 / 62.10 29.35
w/o Turn-Level. 63.35 / 19.65 92.49 / 21.02 41.65 29.60 / 10.85 32.26 / 31.23 22.98
w/o Agent-Specific. 69.55 / 24.55 75.37 / 19.55 40.18 27.00 / 10.50 35.73 / 21.50 22.42
MARSHAL (Hanabi) 71.90 / 7.35 72.52 / 9.29 55.55 26.75 / 5.75 37.36 / 55.12 33.93
w/o Turn-Level. 67.55 / 10.60 68.45 / 31.78 53.20 25.25 / 3.05 54.79 / 47.77 30.68
w/o Agent-Specific. 68.15 / 13.40 74.15 / 10.27 52.50 32.10 / 5.10 44.30 / 56.41 32.08
5 Related work
LLM-based multi-agent systems.
To further extend the ability of LLMs with collaborative intelligence, LLM-based multi-agent systems have been proposed for various tasks. AutoGen (wu2024autogen) and CAMEL (li2023camel) designs cooperative agents for general reasoning and question answering. MetaGPT (hong2024metagpt) and ChatDev (qian2023chatdev) propose static multi-agent workflows with specialized roles for software development tasks. Multi-Agent Debate (liang2023encouraging) takes a competitive approach by letting LLMs propose and criticize solutions over multiple rounds for the final answer. These works mainly focus on designing workflows for multi-agent interactions with fixed LLMs. Other work have proposed to generate workflows automatically (zhang2024aflow). Our work takes a complementary approach by training LLMs to enhance their multi-agent capabilities and build stronger foundation models that can be integrated with these systems.
Reinforcement learning for LLMs.
Reinforcement learning (RL) has emerged as a prominent approach to enhance LLMs’ ability from instruction following to reasoning. To align LLMs with human values, RL from human feedback (RLHF) (ouyang2022training) and from AI feedback (RLAIF) (bai2022constitutional) have been widely used in post-training to steer models towards favorable behaviors. The recent success of reasoning models (jaech2024openai; guo2025deepseek) further popularize RL from verifiable rewards (RLVR) for incentivizing long Chain-of-Thought reasoning capability of LLMs and have been applied in mathematical reasoning (shao2024deepseekmath), coding (wei2025swe), and tool-using (jin2025search). However, these works primarily consider single-turn or single-agent tasks, while our approach extends RL training to multi-turn, multi-agent scenarios. Recently, MT-GRPO (zeng2025reinforcing) also address the turn-level credit assignment issue for multi-turn RL. However, they propose to normalize the turn-level rewards turn-by-turn, which can introduce variance in strategic games with varying number of turns. We propose a "sum-then-normalize" approach that takes advantage of all turns in the normalization, increasing robustness.
Self-play training of LLMs.
The great success of achieving superhuman performance in multi-agent games (silver2016mastering; vinyals2019grandmaster; berner2019dota) has spurred the use of self-play to train LLMs by playing with themselves (zhang2024survey). Some work (chen2024self; wu2024self) combines game-theoretic property of self-play to overcome intransitivity of human preference in LLM alignment. Another line of work (meta2022human; xu2023language; xu2025learning) uses self-play to improve the strategic ability of LLM agents in specific games. The most relevant work to ours is SPAG (cheng2024self) and SPIRAL (liu2025spiral), which also train LLMs via self-play to show generalization to reasoning tasks. However, these work focuses on zero-sum game and show generalization in single-agent evaluations. In contrast, our work consider both cooperative and competitive games and demonstrate generalized improvement in multi-agent system. As a concurrent work to ours, SPIRAL also address the heterogeneity of player role in advantage estimation and introduces a Role-Conditioned Advantage Estimation (RAE), analogous to our agent-specific normalization. We advance this finding by revealing that role separation is only critical in games with distinct return distributions between players.
6 Discussion
In this work, we introduced MARSHAL, a framework utilizing self-play in a diverse range cooperative and competitive strategic games to cultivate multi-agent capabilities in LLMs. Our findings demonstrate that skills honed in strategic games directly transfer to enhanced performance in general multi-agent systems, establishing self-play as a scalable paradigm for training LLM agents.
Despite these promising results, limitations remain. First, our study utilizes two-player games as efficient "prototypes" to cultivate foundational skills. While we demonstrate that these constrained settings are sufficient for strong transfer to general tasks, scaling to larger $N$-player environments introduces significantly greater challenges regarding non-stationarity, population diversity, and credit assignment that warrant dedicated future investigation. Additionally, moving beyond classic games to complex "social sandboxes" (e.g., simulated software engineering, collaborative researching) represents a compelling direction for training agents in more practical domains.
Ultimately, our work provides strong evidence that the principles of self-play are a powerful engine for progress, paving the way for the next generation of sophisticated LLM agents.
Reproducibility statement
All code, model checkpoints, and training scripts required to reproduce the findings of this paper are publicly available at https://github.com/thu-nics/MARSHAL. The repository includes all necessary configurations to replicate our key experiments.
Ethics statement
This work adheres to the ICLR Code of Ethics. The primary ethical consideration of our work is the dual-use nature of enhancing LLM agent capabilities. While our goal is to foster beneficial multi-agent skills like cooperation and competition, the resulting models could potentially be applied to malicious ends. We release our code and models to the research community to encourage further study into the safety and alignment of more capable agents.
Appendix A Use of LLMs
In the preparation of this paper, LLMs were utilized as writing assistants. Their use were limited to improving grammar, clarity, and overall readability. The core research, including the methodology, experiments, and analysis, represents the original work of the authors.
Appendix B Implementation details
Framework and software stack.
Our implementation of MARSHAL is built upon ROLL (wang2025reinforcement), a robust open-source codebase for post-training LLMs with reinforcement learning. ROLL’s native support for agentic, multi-turn rollouts provided a strong foundation for our system. To achieve high performance, ROLL leverages vLLM for efficient inference during the rollout phase (kwon2023efficient) and is built on Megatron-LM for distributed training (megatron-lm). All game environments used in this work were implemented using OpenSpiel (LanctotEtAl2019OpenSpiel) and VS-Bench (xu2025vs), ensuring the correctness and standardization of game logic.
Training settings.
We use Qwen-4B as our base model for all experiments reported in our main results. The training is conducted in a fully online manner, where self-play trajectories are immediately used for policy updates. Specialist agents are trained with a batch size of 128 trajectories, while the generalist agent uses a batch size of 384 (128 from each game). All models were trained for a total of 200 optimization steps. For optimization, we employ the Adam optimizer (kingma2014adam) with a cosine annealing learning rate schedule, which warms up for 10 steps to a peak of $1\times 10^{-6}$ before gradually decaying to 0.
Generation parameters.
During the self-play rollout phase, text is generated via nucleus sampling. We use a temperature of 0.6, a Top-P of 0.99, and a Top-K of 100. These parameters were selected to encourage a diverse yet coherent exploration of strategies.
Hardware configuration.
Models were trained on a single server with 8 NVIDIA H100 GPUs.
Appendix C Evaluation details
Strategic ability.
To rigorously assess game-playing proficiency, we evaluate our agents against a suite of strong, fixed opponents. For competitive games, this includes:
  • Tic-Tac-Toe and Connect Four: two MCTS agents with varying simulation counts (100/1000 for Tic-Tac-Toe, 10/100 for Connect Four) to test against different strengths.
  • Kuhn Poker: The exact Nash Equilibrium (NE) policy.
  • Leduc Hold’em: A NE policy approximated by $5\times 10^{8}$ iterations of Counterfactual Regret Minimization (CFR).
In all competitive games, models are evaluated as both the first-move and second-move player, measured by the average game return. For the cooperative Hanabi variants, performance is measured by the standard self-play game return. In all game settings, agents are evaluated for 1000 games.
For the results on strategic ability reported in the main text, we normalize the scores with respect to the worst-observed performance and the theoretical optimal performance.
For the fixed-opponent ablation studies, the MCTS agent with 100 simulations is used for Tic-Tac-Toe as the opponent, and the NE policy is used for Kuhn Poker.
Generalization to multi-agent systems.
For downstream evaluations, we assess all models on a comprehensive suite of mathematics and general QA benchmarks. In the single-agent setting, we utilize evaluation scripts from Qwen2.5-Math (yang2024qwen2; yang2024qwen2_5_math) for MATH500, GSM8K, AQUA-RAT, AIME24, AMC23, and MMLU-STEM, while employing lm-evaluation-harness (eval-harness) for GPQA-Diamond. For multi-agent evaluations, we adopt the MASLab framework (ye2025maslab) across all benchmarks. To balance evaluation efficiency and accuracy, we set the maximum output length to 8,192 tokens for all settings.
Appendix D Scaling to larger models
To validate the scalability of the MARSHAL framework and ensure our findings are not specific to the 4B parameter scale, we extended our training to the larger Qwen3-8B model. We trained an 8B-parameter MARSHAL generalist agent using the exact same set of strategic games (Tic-Tac-Toe, Kuhn Poker, and Mini Hanabi) and hyperparameters as the 4B experiments.
Strategic ability.
First, we evaluate the model’s proficiency on both training and held-out games. As shown in Table 5, the MARSHAL-8B agent achieves significant improvements over the Qwen3-8B base model across all environments. Notably, we observe the same strong generalization pattern seen in the 4B model: the agent generalizes effectively to held-out, complex games, achieving a drastic improvement in Leduc Hold’em (7.26% to 53.89%) and Simple Hanabi (4.55% to 37.27%).
Table 5: Strategic ability comparison on Qwen3-8B. The MARSHAL training yields consistent improvements on both training games and held-out testing games (Connect Four, Leduc Hold’em, Simple Hanabi).
Model Tic-Tac-Toe Kuhn Poker Mini Hanabi Connect Four Leduc Hold’em Simple Hanabi
Qwen3-8B 48.38 33.12 27.00 10.48 7.26 4.55
MARSHAL (Generalist, 8B) 54.05 44.49 55.28 21.55 53.89 37.27
Generalization to multi-agent systems.
We further evaluate whether these improved strategic capabilities translate to general reasoning benchmarks within multi-agent systems. We integrated the MARSHAL-8B agent into both the competitive MAD framework and the cooperative AutoGen framework.
The results, detailed in Table 6, confirm the scalability of our approach. The MARSHAL-trained 8B model consistently outperforms the base model:
  • In the competitive MAD setting, we observe an average improvement of 2.60%, with substantial gains on challenging math benchmarks like AIME (70.00% to 80.00%).
  • In the cooperative AutoGen setting, the improvement is even more pronounced, with an average gain of 3.90%, and a significant boost on AIME (60.00% to 70.00%).
These results demonstrate that the MARSHAL algorithm and game selection scale stably to larger models, consistently unlocking cooperative and competitive reasoning capabilities.
Table 6: Generalization to multi-agent systems using Qwen3-8B. The MARSHAL generalist consistently improves performance on reasoning benchmarks in both competitive (MAD) and cooperative (AutoGen) settings.
MAS Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MAD Qwen3-8B 82.49 95.00 96.36 83.46 70.00 90.00 89.59 53.03
MARSHAL (Generalist, 8B) 85.09 96.40 96.59 83.46 80.00 95.00 90.70 53.54
AutoGen Qwen3-8B 79.68 88.80 95.91 83.07 60.00 89.19 89.30 51.52
MARSHAL (Generalist, 8B) 83.58 94.40 95.00 85.04 70.00 95.00 90.04 55.56
Appendix E Comparison with SPIRAL: decoupling algorithm and game environments
To rigorously decouple the contributions of our proposed learning algorithm and our selection of game environments, we conducted a set of controlled experiments comparing MARSHAL against SPIRAL baseline methods. We analyze two specific variations:
  1. 1.
    Ablation on algorithm (RAE + our games): We train an agent using SPIRAL’s Role-conditioned Advantage Estimation (RAE) on our full selection of competitive and cooperative game environments. This isolates the impact of the advantage estimation technique.
  2. 2.
    Ablation on games (MARSHAL algorithm + competitive games): We train an agent using the MARSHAL algorithm but restricted to competitive-only games (Tic-Tac-Toe and Kuhn Poker), similar to SPIRAL. This isolates the impact of the cooperative games.
E.1 Strategic ability comparison
Table 7 presents the performance on both training and held-out games. We analyze the results through the lens of the two ablations:
Ablation on algorithm.
When replacing the MARSHAL algorithm with RAE (row 2), we observe a sharp performance decline in the cooperative games. Specifically, performance in Mini Hanabi falls from 54.33% to 24.93%, and in Simple Hanabi from 36.75% to 5.53%. This confirms that RAE struggles with the dense, multi-turn credit assignment required for cooperative settings, validating the necessity of MARSHAL’s turn-level "sum-then-normalize" technique to unlock the potential of self-play learning.
Ablation on games.
When training only on competitive games (row 3), the agent naturally fails to acquire cooperative skills (low scores in Hanabi). This serves as a crucial control baseline for analyzing the downstream transfer results below.
Table 7: Strategic ability comparison decoupling algorithm and game environments. MARSHAL (row 1) outperforms both the RAE-based agent (row 2) and the competitive-only agent (row 3), particularly in cooperative environments.
Model Tic-Tac-Toe Kuhn Poker Mini Hanabi Connect Four Leduc Hold’em Simple Hanabi
MARSHAL (Full Method) 53.13 40.05 54.33 19.98 54.56 36.75
SPIRAL (RAE) + Our Games 48.35 42.05 24.93 9.70 6.11 5.53
MARSHAL Alg. + Adv. Games 48.80 47.55 24.48 17.90 29.98 11.48
E.2 Generalization to multi-agent systems
We further evaluate how these variations affect transfer to general reasoning benchmarks within multi-agent systems.
Generalization to competitive systems (MAD).
As shown in Table 8, the full MARSHAL method achieves the highest average performance (75.96%). Notably, the RAE-based agent underperforms (74.23%), confirming the algorithmic superiority of MARSHAL. The competitive-only agent performs comparably to the full agent, which is expected as MAD is a competitive framework.
Generalization to cooperative systems (AutoGen).
Table 9 reveals the critical role of the game environments. The agent trained on the competitive-only games suffers a significant performance drop in the cooperative AutoGen framework (Average 80.34% vs. 82.15% for MARSHAL). This confirms that cooperative games are not merely "more data"; they are a necessary component for transferring skills to cooperative downstream tasks. Additionally, the RAE-based agent also underperforms, again solidifying the novelty of the algorithmic design of MARSHAL.
Table 8: Generalization to the competitive MAD framework.
Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MARSHAL (Full Method) 75.96 92.80 95.60 83.86 46.67 80.00 87.36 45.45
SPIRAL (RAE) + Our Games 74.23 92.40 96.06 84.25 36.67 77.50 86.79 45.96
MARSHAL Alg. + Adv. Games 75.24 92.20 95.60 83.46 46.67 80.00 86.88 41.92
Table 9: Generalization to the cooperative AutoGen framework. Note the drop in performance for the competitive-only games (row 3), highlighting the necessity of cooperative training.
Model Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MARSHAL (Full Method) 82.15 95.20 94.54 86.61 66.67 92.50 89.53 50.00
SPIRAL (RAE) + Our Games 80.39 94.80 94.77 86.61 60.00 92.50 88.57 45.45
MARSHAL Alg. + Adv. Games 80.34 94.20 94.24 86.61 60.00 90.00 89.37 47.98
Conclusion.
These decoupling experiments confirms the superiority of MARSHAL’s algorithmic design over SPIRAL’s RAE, and that MARSHAL’s successful generalization is not due to one factor alone. It requires both our selection of competitive and cooperative games to provide necessary learning signals, and our novel credit assignment algorithm to effectively learn from them.
Appendix F Comparison with MT-GRPO
A concurrent work MT-GRPO (zeng2025reinforcing) also identifies the limitations of trajectory-level estimation in the original GRPO and proposes a fine-grained turn-level strategy. While MT-GRPO shares our motivation, and is also a critical first step towards fine-grain credit assignment in multi-turn GRPO, our "sum-then-normalize" approach differs fundamentally from their strategy, offering distinct advantages in stability and scalability in the game tasks that we consider in this work.
F.1 Theoretical differences
MT-GRPO relies on normalizing the turn-level rewards turn-by-turn (i.e., computing the mean and variance for the reward at step $t$ across the batch). We identify two structural limitations with this approach that MARSHAL addresses:
  1. 1.
    The variable length problem: In real-world reasoning and gaming tasks, trajectory lengths vary significantly. With turn-by-turn normalization, the effective batch size shrinks as shorter trajectories conclude. This causes the variance of the normalization statistics to increase for later turns in the sequence, introducing instability for training. In contrast, MARSHAL normalizes the cumulative return, which remains defined for the entire trajectory regardless of length.
  2. 2.
    Immediate reward vs. long-term return: MT-GRPO’s local normalization focuses on the immediate step-wise reward relative to peers at that specific step. MARSHAL calculates the full Monte Carlo return first. By normalizing the cumulative outcome, we capture the long-term impact of actions relative to the global baseline, preserving critical long-term dependencies often obscured by local normalization.
F.2 Empirical Comparison
To validate this theoretical analysis, we implemented MT-GRPO and compared it directly against MARSHAL. We trained a Tic-Tac-Toe specialist using both algorithms and evaluated their strategic proficiency across our full game suite.
Strategic ability.
As shown in Table 10, the MT-GRPO agent exhibits notable performance drop across both the training game (Tic-Tac-Toe) and generalization to held-out games (e.g., Connect Four, Hanabi). In particular, the drop is significant in Mini Hanabi (50.48% to 36.67%), a game requiring consistent long-term coordination. This confirms that our "sum-then-normalize" formulation is more robust for the variable-length, multi-turn interactions inherent in strategic games.
Generalization to MAS.
We further extended this comparison to the integration within multi-agent systems to assess if the stability issues in game training affect general reasoning transfer. Table 11 present the results in the competitive MAD and cooperative AutoGen frameworks, respectively. MARSHAL consistently achieves higher average scores (75.01% vs. 73.77% in MAD; 80.15% vs. 79.10% in AutoGen) and outperforms MT-GRPO on key hard reasoning benchmarks like AIME and AMC. This indicates that the training proficiency provided by MARSHAL translates directly to better robust reasoning in general multi-agent tasks.
Table 10: Comparison between MARSHAL and MT-GRPO. The MARSHAL advantage estimation yields superior performance and generalization compared to the concurrent MT-GRPO method, validating the robustness of the "sum-then-normalize" approach.
Model (Tic-Tac-Toe Specialist) Tic-Tac-Toe Kuhn Poker Mini Hanabi Connect Four Leduc Hold’em Simple Hanabi
MARSHAL (Ours) 53.70 38.79 50.48 22.75 43.00 29.75
MT-GRPO 50.10 40.05 36.67 18.55 39.94 20.08
Table 11: Comparison of generalization to multi-agent systems between MARSHAL and MT-GRPO (using Tic-Tac-Toe specialists). MARSHAL achieves higher average performance and demonstrates greater stability in downstream tasks.
MAS Model (Tic-Tac-Toe Specialist) Avg MATH GSM8K AQUA AIME AMC MMLU GPQA
MAD MARSHAL (Ours) 75.01 92.20 96.06 83.07 43.33 82.50 86.76 41.12
MT-GRPO 73.77 92.60 95.91 84.65 36.67 77.50 86.91 42.13
AutoGen MARSHAL (Ours) 80.15 94.40 94.69 87.01 60.00 90.00 89.53 45.45
MT-GRPO 79.10 95.40 94.62 85.04 56.67 90.00 89.02 42.93
Appendix G Hyperparameter analysis on response length penalty
To ensure efficient and stable reasoning within the context window limits, our reward design includes a response length penalty characterized by three hyperparameters: the minimum length $l_{\min}$, the maximum threshold $l_{\max}$, and the penalty weight $\alpha$.
Selection of length constraints.
The length bounds were chosen based on practical engineering constraints:
  • $l_{\min}=11$: This is the hard lower bound determined by the minimum number of tokens required to generate a syntactically valid response structure with the correct format (e.g., <think></think><action>...</action>).
  • $l_{\max}=2048$: This soft upper bound is derived from the context window of the base model. Given Qwen3-4B’s 32k context length, we allocated a budget of approximately 16 turns per game history ($32,768/16=2048$), which provides a sufficient horizon for the reasoning traces in our chosen game portfolio while preventing context overflow.
Ablation on penalty weight $\alpha$.
To assess the sensitivity of the framework to the penalty weight $\alpha$, we conducted an ablation study training Tic-Tac-Toe specialists with $\alpha=1.0$ (stronger regularization), $\alpha=0.5$ (default), and $\alpha=0.0$ (no regularization). The results are detailed in Table 12.
Table 12: Ablation study on the length penalty weight $\alpha$. The default setting ($\alpha=0.5$) strikes a balance between performance and conciseness. Removing the penalty ($\alpha=0$) leads to overly verbose responses and performance degradation in cooperative games like Hanabi. "OL" denotes the percentage of games lost due to overlong responses.
Model Setup Strategic Performance Response Statistics
Tic-Tac-
Toe
Kuhn
Poker
Mini
Hanabi
Connect
Four
Leduc
Hold’em
Simple
Hanabi
Avg Len
OL in
Mini
Hanabi
OL in
Simple
Hanabi
$\alpha=0.5$ (Default) 53.70 38.79 50.48 22.75 43.00 29.75 1700 9.6% 14.3%
$\alpha=1.0$ (Strong) 54.90 42.82 47.53 18.25 50.46 26.05 1657 7.7% 13.9%
$\alpha=0.0$ (None) 54.65 40.95 38.18 23.35 28.61 20.10 1954 20.4% 27.3%
Analysis.
The results yield two key conclusions:
  1. 1.
    Robustness: The framework is robust to reasonable variations in $\alpha$. Increasing the regularization to $\alpha=1.0$ maintains comparable strategic performance to the default setting, indicating that the method is not overly sensitive to the precise magnitude of the penalty.
  2. 2.
    Necessity of regularization: The length penalty is crucial for stability. When removed ($\alpha=0$), the average response length increases significantly (from 1700 to 1954 tokens). This verbosity leads to a sharp performance drop in cooperative, theory-of-mind-demanding games like Mini Hanabi (50.48% to 38.18%). This degradation correlates directly with a dramatic increase in the "Overlong Response" failure rate (9.6% to 20.4%), confirming that the penalty effectively prevents the model from degenerating into computationally wasteful loops.
Appendix H Hyperparameters
To ensure reproducibility and consistency, we maintain a unified hyperparameter configuration for all trained models. The complete settings are provided in Table 13.
Table 13: Hyperparameters
Category Parameter Value
Model Configuration Max Response Length 4096
Max Sequence Length 32768
Dtype bf16
Training Settings Max Steps 200
Train Batch Size 128
Optimizer Adam
Adam parameters ($\beta_{1}$, $\beta_{2}$) (0.9, 0.95)
Learning Scheduler Cosine Annealing
Learning Rate $1\times 10^{-6}$
Weight Decay 0.05
Warm-up Step 10
Gradient Norm Clip 1.0
RL Settings Sampling Temperature 0.5
PPO Epochs 1
(top P, top k) (0.99, 100)
KL Loss True
KL Loss Coefficient 0.2
Entropy Coefficient 0
Dual Clip Loss ture
PPO Policy Clip 0.2
Lambd 0.95
Gamma 1
Appendix I Game observation and prompt
Tic-Tac-Toe
For Tic-Tac-Toe, we provide the agent with a complete observation of the 3x3 game board. The state of each cell—whether it is empty, occupied by ’X’, or occupied by ’O’—is explicitly provided. The prompt clearly indicates which player’s turn it is (’X’ or ’O’) and presents the current board state, asking the agent to select coordinates for its next move from the available empty cells. For example, the game begins with a prompt that provides the empty 3x3 grid and asks the agent to make the first move (Listing 1).
Listing 1: Prompt for Tic-Tac-Toe.
system_prompt:
You are an AI agent that makes optimal decisions to win in the game of Tic-Tac-Toe.
user_prompt:
GAME RULES:
1. Tic-Tac-Toe is a two-player board game played on a three-by-three grid. The grid is 0-indexed, where (0,0) is the top-left corner and (2,2) is the bottom-right corner.
2. Two players take turns placing their marks X and O in empty cells of the grid.
3. The player who first places three of their marks in a horizontal, vertical, or diagonal line wins.
4. If all cells are filled and no player wins, the game ends in a draw.
PLAYER INFORMATION:
1. Your mark is X. You are competing with another player controlling the mark O.
2. In each of your turns:
a. The game state demonstrates the current board with a three-line text grid, where X and O are the marks of the two players, and _ represents empty cells.
b. You need to chose an action to place your mark in an empty cell, based on the given game state and the history of your decisions.
c. All legal actions for the current turn are provided in the format of ‘<X({row},{column})>‘, where X is your mark, and {row} and {column} are integers indicating the row and column of the cell to place your mark.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen action}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><X(0,0)></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
___
___
___
LEGAL ACTIONS:
<X(0,0)>, <X(0,1)>, <X(0,2)>, <X(1,0)>, <X(1,1)>, <X(1,2)>, <X(2,0)>, <X(2,1)>, <X(2,2)>.
Kuhn Poker
For Kuhn Poker, we provide the agent with the observation consists of its single private card (e.g., Jack, Queen, or King) and the complete history actions of two players. The prompt is structured to provide this context clearly, asking the agent to decide on its next action (pass or bet) based on its private information and the betting history. For example, the initial prompt at the beginning of game provides the agent’s private card and the empty action history, asking for the first move (Listing 2).
Listing 2: Prompt for Kuhn Poker.
system_prompt:
You are an AI agent that makes optimal decisions to win in the game of Kuhn Poker.
user_prompt:
GAME RULES:
1. Kuhn Poker is a two-player card game. The deck includes only three cards: King (K) > Queen (Q) > Jack (J).
2. At the start of each game, both player_0 and player_1 place 1 chip into the pot as a blind ante.
3. Each player is dealt a private card, and the third card is set aside unseen.
4. The two players take turns acting, starting with player_0. A player can choose to:
a. <PASS>: place no additional chips into the pot.
b. <BET>: place 1 additional chip into the pot.
5. If a player chooses to <PASS> after the other players <BET>, the betting player wins the pot.
6. If both players choose to <PASS> or both players choose to <BET>, the player with the higher card wins the pot.
PLAYER INFORMATION:
1. You are player_0. You are competing with player_1.
2. In each of your turns:
a. The game state shows your private card and the betting history.
b. You need to choose an action based on your card and the current game state.
c. All legal actions for the current turn are provided in the format of ‘<PASS>‘ or ‘<BET>‘.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen action}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><PASS></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
1. Blind ante: both player_0 and player_1 place 1 chip into the pot.
2. Deal: your card is Jack (J).
LEGAL ACTIONS:
<PASS>, <BET>.
Hanabi
For the Hanabi variants, a cooperative imperfect information game, we provide the agent with a unique observation. An agent observes the cards held by all other players but remains unaware of its own hand. The observation also includes the number of remaining information and life tokens, the cards successfully played on the board (fireworks), and the contents of the discard pile. The prompt is highly structured, detailing the game rules, player-specific information, and strict response formatting. For training, we use Mini Hanabi, a simplified 2-player version with 2 colors, 2 ranks, 3-card hands, and 3 information/life tokens. For out-of-distribution evaluation, we use a more complex variant, Simple Hanabi, which features 3 colors, 2 ranks, 5-card hands, 8 information tokens and 3 life tokens. For example, an initial prompt details the starting setup of the chosen variant, including all visible cards and token counts, before asking the first player to make the first move (Listing 3).
Listing 3: Prompt for Mini Hanabi.
system_prompt:
You are an AI agent that makes optimal decisions to achieve the highest score in the game of Hanabi.
user_prompt:
GAME RULES:
1. Hanabi is a cooperative card game for 2 players, player 0 and player 1.
2. The deck consists of 2 colors: Red(denoted by R), Yellow(denoted by Y), with ranks ranging from 1 to 2. Each color contains 4 cards: three of rank 1, and one of rank 2, for a total of 8 cards.
3. Each player holds 3 cards in hand. Players can observe the hand of the other player, but not their own.
4. There are 3 information tokens and 3 life tokens shared by both players.
5. The objective is to play cards in ascending order of rank, from 1 to 2, to their corresponding color stacks, hence achieving the Fireworks’.
6. The players take turns to take one of the following actions:
a. <Play i‘>: play the i-th card from the players own hand (0-indexed). If the card is sequential to the top card of its corresponding color stack, the move is valid and the card is added to the top of the stack, then both players receive 1 point. Otherwise, a life token is lost.
b. <Discard i‘>: discard the i-th card from the players own hand and gain one information token.
c. <Reveal player +1 color c‘>: spend one information token to reveal all cards of color c in the other players hand.
d. <Reveal player +1 rank r‘>: spend one information token to reveal all cards of rank r in the other players hand.
7. After playing or discarding, the player receives a new card from the deck (if remaining).
8. The game ends when:
a. If all color stacks are completed (i.e., all cards of rank 2 are played to their corresponding color stacks), then both players finish the game with the highest possible total score of 4.
b. If deck is depleted, both players finish the game with a total score which equals the sum of the highest ranks of each color stack.
c. If all life tokens are lost before the above two conditions are met, then both players lose all points they have earned so far, and finish the game with a total score of 0.
PLAYER INFORMATION:
1. You will be playing as the player 0.
2. In each of your turns, you will be provided with the current game state information, including the remaining life tokens and information tokens, the current color stacks, the remaining deck size, the discard pile, the hand of the other player, and the revealed information on your own hand.
3. Known cards are denoted by their color and rank. For example, R2 means a red card of rank 2. 4. The current color stacks are represented by the top card of each color stack. In particular, rank 0 denotes an empty stack. For example, Y0 means the yellow stack is still empty.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen action}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><Discard Card 0></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of all life tokens and end of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
1. There are 3 life tokens and 3 information tokens remaining.
2. The top of the color stacks are: R0 Y0.
3. 2 cards remain in the draw pile.
4. The discard pile currently contains: None.
5. The other players hand:
- Card 0 (Y1): the other player believes it is one of the colors [R, Y] and one of the ranks [1, 2].
- Card 1 (R1): the other player believes it is one of the colors [R, Y] and one of the ranks [1, 2].
- Card 2 (R1): the other player believes it is one of the colors [R, Y] and one of the ranks [1, 2].
6. Your own hand, based on the revealed information:
- Card 0: one of the colors [R, Y] and one of the ranks [1, 2].
- Card 1: one of the colors [R, Y] and one of the ranks [1, 2].
- Card 2: one of the colors [R, Y] and one of the ranks [1, 2].
LEGAL ACTIONS:
<Play card 0>, <Play card 1>, <Play card 2>, <Reveal player +1 color R>, <Reveal player +1 color Y>, <Reveal player +1 rank 1>.
Connect Four
For Connect Four, we provide the agent with a complete observation of the 6x7 game board, showing the positions of all ’X’ and ’O’ marks. The prompt is designed to be comprehensive, initially presenting the rules of the game (connecting four pieces to win, draw conditions). It then informs the agent of its assigned mark (’X’ or ’O’) and the opponent’s mark. For each turn, the prompt presents the current board state and asks the agent to choose a column (0-6) to drop its piece. For example, the game begins with a prompt that provides the empty 6x7 grid, explains the rules and player marks, and asks the first player to choose a column (Listing 4).
Listing 4: Prompt for Connect Four.
system_prompt:
You are an AI agent that makes optimal decisions to win in the game of Connect Four.
user_prompt:
GAME RULES:
1. Connect Four is a two-player board game played on a 6x7 grid. Players take turns dropping their pieces into columns.
2. The goal is to connect four of your pieces horizontally, vertically, or diagonally.
3. Pieces fall to the bottom of the column or stack on top of existing pieces.
4. The first player to connect four pieces wins. If the board fills up without a winner, its a draw.
PLAYER INFORMATION:
1. Your mark is X. You are competing with another player controlling the mark O.
2. In each of your turns:
a. The game state shows the current board as a 6x7 grid.
b. You need to choose a column (0-6) to drop your piece, where 0 denotes the leftmost column, 6 denotes the rightmost column.
c. All legal actions are provided as ‘<X({column})>‘, where X is your mark, and {column} is the column number.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen column}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><X(3)></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
_______
_______
_______
_______
_______
_______
LEGAL ACTIONS:
<X(0)>, <X(1)>, <X(2)>, <X(3)>, <X(4)>, <X(5)>, <X(6)>.
Leduc Hold’em
For Leduc Hold’em, we provide the agent with the observation which contains all necessary information for decision-making. The observation is more complex due to two betting rounds and a public card. In the first round, the agent observes its private card and the betting history. In the second round, after a public card is revealed, the observation is updated to include this board card. The prompt provides the agent with its private card, the current betting history, and the public card if revealed, requiring it to make an action in the context of the evolving game state. For example, the initial prompt at the beginning of game provides the agent with its private card and asks for an action in the first betting round before any actions have been taken (Listing 5).
Listing 5: Prompt for Leduc Hold’em.
system_prompt:
You are an AI agent that makes optimal decisions to win in the game of Leduc Poker.
user_prompt:
GAME RULES:
1. Leduc poker is a two-player card game. The deck includes only six cards: two pairs of King (K), Queen (Q), and Jack (J).
2. At the start of each game, both player_0 and player_1 place 1 chip into the pot as a blind ante.
3. Each player is dealt one private card from the deck, and the remaining cards are set aside unseen.
4. The game has two betting rounds. When the first round ends, one public card from the remaining cards of the deck is revealed to both players.
5. The two players take turns acting in the betting rounds, both starting with player_0. A player can choose to:
a. <FOLD>: stop betting and the other player wins the pot.
b. <CALL>: match the current bet. If no bet has been made in the current round, this is equivalent to checking.
c. <RAISE>: first match the current bet and then add n chips to the bet, where n=2‘ in the first round and n=4‘ in the second round. If no bet has been made in the current round, this is equivalent to betting n chips.
6. A maximum of two <RAISE>s are allowed in each round. Each round ends when both players have acted and their bets are equal.
7. If a player chooses to <FOLD>, the other player wins the pot.
8. If neither player chooses to <FOLD>, the second round ends with a showdown:
a. If a player has a pair (private card = public card), the player wins the pot.
b. If neither player has a pair, the player with the higher card (K > Q > J) wins the pot.
c. If two players have the same card, the players split the pot.
PLAYER INFORMATION:
1. You are player_0. You are competing with player_1.
2. In each of your turns:
a. The game state shows your private card, public card (if revealed), and the betting history.
b. You need to choose an action based on your cards and the current game state.
c. All legal actions for the current turn are provided in the format of ‘<FOLD>‘, ‘<CALL>‘, or ‘<RAISE>‘.
RESPONSE INSTRUCTIONS:
Always choose only one action from the legal actions and output ‘<answer>{your chosen action}</answer>‘ with no extra text after you finish the thinking process. For example, ‘<answer><CALL></answer>‘. Strictly follow the above format and keep your thinking process concise. Responses that do not follow the format will result in immediate loss of the game.
Information of Turn-1:
This is your turn. The game state and legal actions for this turn are provided below. Please choose your action and strictly follow the given output format in the response instructions.
GAME STATE:
1. Blind ante: both player_0 and player_1 place 1 chip into the pot.
2. Deal: your card is J.
LEGAL ACTIONS:
<CALL>, <RAISE>.
Appendix J Full results
In this section, we present the raw (unnormalized) results for all experiments in the main text.
Table 14: Full results of game-play return in training games
Model Tic-Tac-Toe Kuhn Poker Mini Hanabi
MCTS
(100 Simulations)
MCTS
(1000 Simulations)
Qwen3-4B 0.403/-0.629 -0.374/-0.687 -0.148/-0.141 1.200
SPIRAL 0.470/-0.624 -0.297/-0.717 -0.301/-0.149 1.494
MARSHAL
   Tic-Tac-Toe 0.506/-0.358 -0.278/-0.379 -0.119/-0.142 2.019
   Kuhn Poker 0.397/-0.490 -0.388/-0.538 -0.107/-0.103 1.799
   Mini Hanabi 0.438/-0.644 -0.350/-0.702 -0.123/-0.130 2.222
   Generalist 0.544/-0.419 -0.302/-0.450 -0.114/-0.141 2.173
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe 0.760/-0.161 -0.378/-0.162 -0.146/-0.090 1.397
   Kuhn Poker -1.000/-1.000 -1.000/-1.000 -0.114/-0.117 0.000
   Mini Hanabi - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe 0.492/-0.517 -0.292/-0.567 -0.104/-0.091 1.392
   Kuhn Poker 0.267/-0.607 -0.426/-0.625 -0.074/-0.106 1.666
   Mini Hanabi 0.351/-0.788 -0.400/-0.811 -0.133/-0.084 2.128
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe 0.654/-0.376 -0.213/-0.366 -0.127/-0.126 1.764
   Kuhn Poker 0.391/-0.509 -0.446/-0.567 -0.116/-0.109 1.607
   Mini Hanabi 0.363/-0.732 -0.364/-0.764 -0.119/-0.128 2.100
Table 15: Full results on game-play return in testing games
Model Connect Four Leduc Hold’em Simple Hanabi
MCTS
(10 Simulations)
MCTS
(100 Simulations)
Qwen3-4B -0.462/-0.912 -0.902/-0.998 -0.629/-0.691 0.833
SPIRAL -0.383/-0.833 -0.868/-0.995 -0.870/-0.706 0.836
MARSHAL
   Tic-Tac-Toe -0.387/-0.703 -0.858/-0.974 -0.518/-0.702 1.785
   Kuhn Poker -0.448/-0.746 -0.874/-0.995 -0.460/-0.327 1.761
   Mini Hanabi -0.465/-0.885 -0.880/-0.998 -0.736/-0.403 2.036
   Generalist -0.421/-0.780 -0.871/-0.979 -0.483/-0.487 2.205
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe -0.593/-0.887 -0.939/-0.990 -0.632/-0.616 0.733
   Kuhn Poker -1.000/-1.000 -1.000/-1.000 -1.124/-1.003 0.000
   Mini Hanabi - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe -0.465/-0.754 -0.843/-0.989 -0.622/-0.553 1.143
   Kuhn Poker -0.408/-0.783 -0.899/-0.998 -0.789/-0.663 1.379
   Mini Hanabi -0.495/-0.939 -0.886/-0.998 -0.555/-0.483 1.841
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe -0.492/-0.790 -0.882/-0.983 -0.594/-0.460 1.303
   Kuhn Poker -0.460/-0.790 -0.908/-0.992 -0.753/-0.769 1.345
   Mini Hanabi -0.358/-0.898 -0.902/-0.998 -0.664/-0.389 1.925
Table 16: Full results on general benchmarks in standard single agent setting
Model MATH GSM8K AQUA AIME AMC MMLU GPQA
Qwen3-4B 87.60 94.60 39.80 36.70 70.00 57.10 39.39
SPIRAL 87.50 94.80 51.20 36.70 80.00 58.70 37.37
MARSHAL
   Tic-Tac-Toe 89.10 95.20 46.50 40.00 77.50 57.60 38.89
   Kuhn Poker 87.80 94.50 48.40 33.30 72.50 59.30 33.84
   Mini Hanabi 88.10 94.70 48.00 43.30 65.00 58.90 36.36
   Generalist 89.90 94.60 52.00 33.30 75.00 59.90 34.85
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe 89.10 94.80 53.90 50.00 72.50 63.00 37.88
   Kuhn Poker 86.20 94.00 45.70 33.30 72.50 53.80 33.84
   Mini Hanabi - - - - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe 88.80 94.80 48.80 40.00 72.50 57.70 36.36
   Kuhn Poker 88.60 94.00 48.40 40.00 72.50 57.90 34.34
   Mini Hanabi 87.80 94.50 48.40 36.70 70.00 59.30 36.87
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe 88.20 94.50 53.10 36.70 77.50 58.20 39.39
   Kuhn Poker 88.10 94.80 53.90 40.00 75.00 58.70 35.86
   Mini Hanabi 88.60 93.90 46.10 40.00 75.00 57.00 38.89
Table 17: Full results on general benchmarks using the MAD framework
Model MATH GSM8K AQUA AIME AMC MMLU GPQA
Qwen3-4B 90.20 95.91 80.71 40.00 75.00 87.42 37.88
SPIRAL 91.60 95.45 81.89 40.00 77.50 87.01 40.40
MARSHAL
   Tic-Tac-Toe 92.20 96.06 83.07 43.33 82.50 86.76 41.12
   Kuhn Poker 91.60 96.21 82.68 40.00 82.50 87.39 41.41
   Mini Hanabi 91.40 95.60 82.68 43.33 77.50 87.04 38.38
   Generalist 92.80 95.60 83.86 46.67 80.00 87.36 45.45
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe 94.80 94.54 85.04 53.33 90.00 88.57 42.93
   Kuhn Poker 95.00 94.24 86.22 53.33 85.00 88.54 45.45
   Mini Hanabi - - - - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe 93.80 94.77 85.04 60.00 85.00 88.98 47.98
   Kuhn Poker 95.00 94.54 86.22 56.67 90.00 89.05 50.00
   Mini Hanabi 94.20 94.62 85.43 53.33 90.00 88.51 46.97
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe 94.40 94.62 84.65 56.67 87.50 89.33 46.97
   Kuhn Poker 95.00 94.69 84.25 56.67 85.00 89.62 47.47
   Mini Hanabi 94.20 94.47 87.01 63.33 90.00 88.92 43.94
Table 18: Full results on general benchmarks using the AutoGen framework
Model MATH GSM8K AQUA AIME AMC MMLU GPQA
Qwen3-4B 93.40 94.69 85.04 56.67 87.50 89.21 47.47
SPIRAL 94.20 94.47 86.61 60.00 87.50 91.60 45.96
MARSHAL
   Tic-Tac-Toe 94.40 94.69 87.01 60.00 90.00 89.53 45.45
   Kuhn Poker 95.80 94.39 86.61 63.33 92.50 89.65 48.48
   Mini Hanabi 94.40 94.54 86.22 66.67 95.00 88.98 44.95
   Generalist 95.20 94.54 86.61 66.67 92.50 89.53 50.00
MARSHAL (w/ Fixed Opponent)
   Tic-Tac-Toe 92.40 95.38 83.07 43.33 80.00 87.11 43.94
   Kuhn Poker 90.20 95.83 83.46 33.33 72.50 86.69 42.42
   Mini Hanabi - - - - - - -
MARSHAL (w/o Turn-Level Advantage Estimator)
   Tic-Tac-Toe 91.00 95.75 82.28 40.00 72.50 86.34 42.93
   Kuhn Poker 91.80 95.45 82.68 36.67 82.50 86.79 41.92
   Mini Hanabi 91.20 95.91 83.86 40.00 75.00 87.33 41.92
MARSHAL (w/o Agent-Specific Advantage Normalization)
   Tic-Tac-Toe 91.20 95.83 83.46 40.00 80.00 87.14 38.58
   Kuhn Poker 91.20 95.53 84.65 36.67 75.00 86.63 43.94
   Mini Hanabi 92.60 96.06 82.68 33.33 80.00 87.11 43.43